### v1

# Not all Chess960 positions are equally complex

Abstract
We analyze strategic complexity across all 960 Chess960 (Fischer Random Chess) starting positions. Stockfish evaluations show a near-universal first-move advantage for White ($\langle E\rangle=+0.30\pm 0.14$ pawns), indicating that the advantage conferred by moving first is a robust structural feature of the game. To quantify decision difficulty, we introduce an information-based measure $S(n)$ describing the cumulative information required to identify optimal moves over the first $n$ plies. This measure decomposes into contributions from White and Black, $S_{W}$ and $S_{B}$, yielding a total opening complexity $S_{\mathrm{tot}}=S_{W}+S_{B}$ and a decision asymmetry $A=S_{B}-S_{W}$. Across the ensemble, $S_{\mathrm{tot}}$ varies by a factor of three, while $A$ spans from 2.5-2.5 to +1.8+1.8 bits, showing that some openings burden White and others Black. The mean $\langle A\rangle=-0.25$ bits indicates a slight tendency for White to face harder opening decisions. Standard chess (position #518, RNBQKBNR) exhibits above-average asymmetry (91st percentile) but typical overall complexity (47th percentile). The most complex opening is #226 (BNRQKBNR), whereas #198 (QNBRKBNR)is the most balanced, with both evaluation and asymmetry near zero. These results reveal a highly heterogeneous Chess960 landscape in which small rearrangements of the back-rank pieces can significantly alter strategic depth and competitive fairness. Remarkably, the classical starting position—despite centuries of cultural selection—lies far from the most balanced configuration.
I Introduction
Chess, one of humanity’s oldest and most studied strategic games, served as a simple testing ground for theories of decision-making, artificial intelligence, and complex systems [1]. From a physics perspective, chess represents an ideal model system for studying decision-making complexity and information dynamics: it is fully deterministic, has well-defined rules, generates vast empirical datasets, and admits precise computational analysis of optimal strategies. The modern rules of chess crystallized in the 15th century, with the classical starting position—RNBQKBNR (Rook-Knight-Bishop-Queen-King-Bishop-Knight-Rook)—becoming the universal standard. This position was not derived from mathematical principles but emerged through cultural evolution and practical play over centuries [2].
The dominance of opening theory in modern chess has led to a paradoxical situation: at the highest levels of play, the first 15–20 moves often follow extensively analyzed variations stored in databases containing millions of games. This ‘opening book’ knowledge can extend so deeply that the creative, analytical phase of the game is delayed until the middlegame. The tension between memorization and genuine understanding has long concerned chess theorists. As early as 1792, the Dutch chess enthusiast Philip Julius van Zuylen van Nijevelt proposed randomizing starting positions to restore creative play [3]. Two centuries later, World Champion Bobby Fischer recognized the same issue as a potential threat to the game’s intellectual vitality [4]. In 1996, Fischer proposed a variant that preserves chess’s strategic depth while eliminating memorization advantages: Fischer Random Chess, developed in collaboration with Susan Polgar [4]. Now standardized as Chess960 or ‘Fischer random’, and more recently promoted as ‘Freestyle Chess’ [5], this variant shuffles the pieces on the back rank subject to three constraints: (i) the two bishops must occupy opposite-colored squares, (ii) the king must be positioned between the two rooks to preserve castling rights, and (iii) White and Black must share identical mirrored arrangements. These rules generate exactly $960$ legal starting positions, with the classical arrangement RNBQKBNR corresponding to index #518 in the standard numbering.
The number $960$ follows directly from elementary combinatorics. Labeling the back-rank squares (from a to h), the bishops can be placed on any pair of light and dark squares, giving $4\times 4=16$ combinations. The queen may then occupy any of the remaining six squares. This leaves five squares to be filled by the multiset $\{R,K,R,N,N\}$. With 5 squares left, there are $\binom{5}{2}=10$ ways to place the two knights; the remaining three squares must then be occupied by the two rooks and the king, with the king between the rooks, which yields a unique arrangement for these three pieces. Altogether, this gives $16\times 6\times 10=960$ possible starting positions. This deterministic construction also provides a canonical indexing scheme from 0 to $959$, known as the Scharnagl enumeration algorithm [6, 7]. In this scheme, each starting position is uniquely encoded by three integers: (i) the bishop code, selecting one of the $16$ opposite-colored bishop pairs; (ii) the queen position, selecting one of the $6$ remaining squares after the bishops are placed; and (iii) the N5N code, indexing one of the admissible arrangements of the multiset $\{R,K,R,N,N\}$ on the final five squares under the requirement that the king lies between the two rooks. The mapping to an index is given by $\mathrm{idn}\;=\;(\text{bishop code})\;+\;16\times(\text{queen position})\;+\;96\times(\text{N5N code})$. For the classical starting position RNBQKBNR, this yields $\mathrm{idn}=6\;+\;16\times 2\;+\;96\times 5=518$ which is the standard numbering for classical chess. This indexing provides a compact and reproducible method for scanning all $960$ starting configurations and guarantees a one-to-one correspondence between integers in $[0,959]$ and legal Chess960 starting positions.
The variant gained official recognition when FIDE incorporated it into the Laws of Chess in 2008 and began sanctioning world championships in 2019 [4]. More recently, freestyle chess gained more attention with the new ‘Freestyle Chess Grand Slam Tour’ [5]. From the standpoint of statistical physics, Chess960 offers a unique opportunity: a discrete ensemble of 960 initial conditions evolving under identical dynamical rules, enabling systematic comparison of complexity across configurations. While Fischer’s motivation was practical—restoring the importance of understanding over memorization—his innovation raises questions amenable to quantitative analysis: Are all 960 positions equally complex? Does the classical position occupy a special place in the complexity landscape? Can we quantify the ‘decision-making difficulty’ intrinsic to each starting configuration? These questions parallel fundamental inquiries in statistical physics concerning how initial conditions influence the complexity of dynamical trajectories.
Recent advances in computer chess, particularly the development of engines like Stockfish capable of evaluating positions with superhuman accuracy [8], enable us to address these questions rigorously. The availability of large databases has enabled the emergence of a quantitative ‘science of chess,’ with studies addressing innovation dynamics [9], statistical regularities in opening frequencies [10], network properties of move sequences [11, 12, 13, 14, 15], memory effects and temporal correlations [16, 17], game complexity measures [18, 19, 20], human performance and cognition [21, 22, 23], and spatial control patterns [24, 25]. These works demonstrate that chess provides a rich empirical system for studying decision-making under uncertainty.
In this work, we characterize the 960 initial positions using simple quantitative measures, and we introduce an information–cost measure $S(n)$ that captures the cumulative information required to identify optimal moves over the first $n$ plies of gameplay. We apply this measure systematically to the 960 Chess960 positions, including the classical position #518, to construct an empirical ‘complexity landscape’ of opening configurations. Our approach treats chess as a complex system where each move represents a decision under uncertainty, with $S(n)$ capturing the difficulty of this decision-making process. This framework allows us to investigate whether the classical starting position—selected by historical accident rather than optimization—occupies any privileged location in the space of possible chess games, and to identify configurations that may offer greater symmetry or fairness between White and Black.
II Results
II.1 Simple measures
II.1.1 Location probability
Figure 1: Probability distribution of pieces for the 960 initial positions across back-rank squares (from a to h). The bishops, knights, and queen each occupy any of the eight back-rank squares with equal probability $1/8$. By contrast, the king is more likely to appear near the central files, while the rooks are correspondingly more likely to be placed toward the sides.
As discussed above, there are 960 possibilities that result from the following constraints: The two bishops must occupy opposite-colored squares (one on a light square, one on a dark square); the king must be positioned between the two rooks, preserving the ability to castle kingside and queenside. Finally, we note that the Black pieces mirror the White arrangement, ensuring a priori fairness. These 960 positions are numbered from 0 to 959 and the standard chess position (RNBQKBNR) corresponds to the number 518 in this numbering [26]. The simplest question we can ask under these constraints is: what is the probability that a given piece occupies a particular file? We show this probability in Fig. 1. The bishops, knights, and the queen have equal probability across all eight squares ($1/8=12.5\%$). Due to castling constraints, the king and the rooks have non-uniform probability distributions across the eight squares. The king cannot be on the a or h file and we observe that it is most likely close to the central files (d-e), and naturally, the rooks have a larger probability to be on the sides. Fig. 1 shows the probability distribution for each piece type across the eight squares, revealing the symmetry and constraints of the Chess960 rules.
II.1.2 Initial Stockfish evaluation
To quantify the structural asymmetry inherent in Chess960’s turn-based mechanics, we evaluated all 960 starting positions using Stockfish 17.1 [8] at depth 30. The resulting distribution (Fig. 2) reveals a striking uniformity: the mean initial evaluation is $\langle E\rangle=+0.297\pm 0.136$ pawns, with 956 of 960 positions (99.6%) exhibiting positive evaluation favoring White. Only four position achieve exact or close to balance. This near-universal first-move advantage confirms that the privilege of moving first confers a measurable strategic benefit independent of piece configuration.
Standard chess (position #518, RNBQKBNR) exhibits an initial evaluation of +0.30+0.30 pawns, placing it in the 47th percentile. This demonstrates that the classical starting position is statistically typical within the Chess960 landscape, neither amplifying nor diminishing the structural benefit of initiative.
Figure 2: Distribution of Stockfish initial position evaluations across all 960 Chess960 starting positions (Stockfish 17.1, depth 30). The distribution is centered at $\langle E\rangle=+0.30\pm 0.14$ pawns, with 99.6% of positions favoring White. Vertical lines indicate position #518 (standard chess, solid green), position #279 (maximum White advantage, red dashed), position #535 (almost perfect balance, blue dashed), and the ensemble mean (black dashed).
The distribution’s relatively narrow standard deviation ($\sigma=0.14$ pawns) indicates remarkable consistency across disparate piece arrangements, suggesting that White’s first-move advantage is a robust structural feature largely independent of the specific configuration. The most White-favorable position (#279, NRBKNRQB) reaches +0.83+0.83 pawns—approximately 2.8 times the mean. This configuration places in particular the rook, the queen and a bishop on the f-h files, creating immediate developmental asymmetries that White can exploit (although the engine’s sequence of best moves starting with f4, c4, Nf3 leads to variations that are not obvious). Notably, position #535 (RNBKQNRB) achieves almost perfect balance (with an infinitesimal advantage for Black at 0.09-0.09 pawns) —one of only three configurations where the first-move advantage is close to null. The rarity of balanced or Black-favoring configurations (0.4% of positions) underscores that neutralizing White’s first-move advantage requires a precise alignment of spatial and tactical factors. Note that these results depend on the engine and search depth. While specific outliers may shift, the overall distribution and its statistics remain robust.
From a game-theoretic perspective, the universality of White’s advantage across Chess960 positions suggests that the first-move privilege is not an artifact of centuries of opening theory refinement in standard chess, but rather an intrinsic property of the game’s mechanics. The advantage magnitude of $\approx 0.3$ pawns—roughly one-third of a minor piece—is consistent with empirical win-rate statistics from high-level play, where White scores a little above 50% in standard chess. Fischer’s motivation for Chess960 was to eliminate memorized opening theory, not to eliminate the first-move advantage itself, and our results confirm that this asymmetry persists as a fundamental feature of the game independent of initial configuration. This finding has implications for tournament fairness: while Chess960 successfully neutralizes preparation advantages, it preserves the need for color-balanced pairings or double-game formats to ensure equity between opponents.
II.1.3 Branching factor and position space growth analysis
To quantify the complexity evolution across different Chess960 starting positions, we analyze two key metrics: the branching factor $b(n)$ and the position space size $N(n)$ (i.e. the number of different positions) as functions of ply depth $n$. This analysis characterizes the raw combinatorial complexity of the game tree, independent of strategic considerations or optimal play. It quantifies the fundamental decision-making burden at each stage of the game: how many distinct game states a player must consider when planning ahead.
For each position, we compute $N(n)$ defined as the number of unique legal positions reachable at ply depth $n$. We then compute the branching factor $b(n)$ defined as the ratio of consecutive position counts
$b(n)=\frac{N(n+1)}{N(n)}$ (1)
This ratio quantifies the average number of accessible successor states per position at depth $n$. Figure 3 presents a comparative analysis between position 518 (standard chess) and the ensemble average across all sampled positions. Panel (a) shows the evolution of the branching factor $b(n)$ versus ply depth. The average branching factor across sampled positions (blue curve with error bars representing one standard deviation) exhibits typical values of order $20$ consistent with previous estimates [24]. Note that branching factors around $b\sim 35$, often cited in computer chess literature, correspond to middlegame positions where piece development leads to increased move options [24]. Position #518 (green curve) closely tracks this average, with deviations remaining within one standard deviation throughout the analyzed range.
Panel (b) displays the growth of unique positions $N(n)$ on a semi-logarithmic scale. The approximately linear behavior confirms exponential growth, $N(n)\sim\exp(\lambda n)$. The extremely small dispersion around the average (smaller than the symbol size) indicates that all starting positions generate state spaces of essentially identical size. Note that the effective growth rate $\lambda$ is notably smaller than what would be predicted from the branching factor alone ($\lambda<\ln\langle b\rangle$). This reduction arises because different move sequences can converge to the same board configuration—a phenomenon known as transposition in chess terminology—causing the position space to grow more slowly than the space of legal move sequences.
Figure 3: (a) Average branching factor $\langle b\rangle$ versus ply depth $n$ across all 960 starting configurations (blue), compared with standard chess (position #518, green). Error bars indicate one standard deviation. The dashed line marks the commonly cited middlegame value $b\approx 35$. (b) Number of unique positions $N(n)$ versus ply depth on a semi-logarithmic scale, showing exponential growth $N(n)\sim b_{\mathrm{eff}}^{n}$ with effective branching factor $b_{\mathrm{eff}}\approx 14.41$, significantly smaller than $\langle b\rangle$ due to transpositions.
From the perspective of branching factors and the number of unique positions, our results show no significant dependence on the initial arrangement; in this sense, all 960 starting positions are equivalent.
II.2 Characterizing the opening complexity
II.2.1 Asymmetry in Decision Complexity
The measures discussed so far, however, do not quantify how difficult it is to play a given position as White or Black, nor do they assess whether one starting configuration is intrinsically more complex than another. To address these questions, we introduce an information-based measure of decision complexity that captures the difficulty of identifying the best moves.
Consider a position where a chess engine evaluates (in centipawns) the best move with score $E_{1}$ and the second-best move with score $E_{2}$. The difference $\Delta=E_{1}-E_{2}$ quantifies how much ‘better’ the optimal move is [24, 25]. For a player with discrimination ability $\Delta_{0}$ (the minimum evaluation difference they can reliably perceive), we model the probability of choosing the optimal move (over the second best move) using a Boltzmann-like factor (also called the softmax function)
$P(\text{optimal})=\frac{1}{1+e^{-\Delta/\Delta_{0}}}$ (2)
This form has the expected limiting behavior: when $\Delta\gg\Delta_{0}$, $P\to 1$ (the best move is obvious), and when $\Delta\ll\Delta_{0}$, $P\to 1/2$ (almost indistinguishable moves). To quantify the difficulty of identifying the optimal move at a given ply, we associate to each position an information content or cost (in bits) given by
$S(\Delta)=-\log_{2}P=\log_{2}\!\left(1+e^{-\Delta/\Delta_{0}}\right),$ (3)
where $\Delta$ is the evaluation gap between the best and second-best moves, and $\Delta_{0}$ sets the discrimination scale (analogous to a noise or ‘temperature’ parameter). Importantly, $S(\Delta)$ is not an entropy of a probability distribution. Instead, it measures the amount of information required to resolve the local ambiguity in the game tree. The form (3) is identical to the log-partition function of a two-state softmax (logit) choice model, widely used in decision theory and statistical physics. In that context, $S(\Delta)$ plays the role of the ‘cost’ of discriminating between nearly equivalent alternatives. This interpretation yields intuitive limits:
  • If $\Delta\ll\Delta_{0}$, the two moves are nearly indistinguishable: identifying the optimal one requires close to a full bit of information, $S(\Delta)\approx\log_{2}2=1$ bit.
  • If $\Delta\gg\Delta_{0}$, the best move is effectively forced: the information cost collapses to zero, $S(\Delta)\approx 0$ bit.
For a trajectory of $n$ plies, the cumulative information required to navigate the resulting branch of the search tree is
$S(n)=\sum_{i=1}^{n}S(\Delta_{i})=\sum_{i=1}^{n}\log_{2}\!\left(1+e^{-\Delta_{i}/\Delta_{0}}\right),$ (4)
with $\Delta_{i}$ the evaluation gap at ply $i$. The average information per move, $S(n)/n$, reflects the typical decision difficulty encountered along the trajectory. $S(n)$ provides a principled, information-theoretic measure of how demanding a given sequence of moves (such as an opening for example) is to navigate under optimal play.
The quantity $S(n)$ has several natural properties: it is monotonically increasing with $n$ (information accumulates along the line of play); forced positions with large $\Delta_{i}$ contribute negligibly; subtle positions with small $\Delta_{i}$ dominate; and the overall magnitude depends on playing strength through $\Delta_{0}$. Typical values are $\Delta_{0}\approx 10$ cp for Grandmaster level, $\Delta_{0}\approx 50$ cp for Expert level, and $\Delta_{0}\approx 100$ cp for intermediate level [25]. We focus on the Grandmaster level here and set $\Delta_{0}=10$ cp.
[…]
Figure 4: Information cost analysis across the 960 Chess960 starting positions based on $n=10$ plies per player of optimal play (Stockfish 17.1, depth 14, discrimination threshold $\Delta_{0}=10$ cp). Values represent averages over 5 games per position. (a) Distribution of White’s cumulative cost $S_{W}(10)$ showing the decision complexity faced by White across different starting configurations. The green vertical line marks position #518 (standard chess) at $\langle S_{W}\rangle=4.29$ bits, near the ensemble mean. (b) Distribution of Black’s cumulative cost $S_{B}(10)$. Position #518 exhibits $\langle S_{B}\rangle=4.97$ bits, slightly above the ensemble mean. (c) Asymmetry $A=S_{B}-S_{W}$ versus total information cost $S_{\mathrm{tot}}=S_{W}+S_{B}$. Negative asymmetry indicates White faces more complex decisions; positive values indicate Black faces greater complexity. Standard chess (position #518, indicated by a gold star) shows positive asymmetry ($\langle A\rangle=+0.69$ bits) and moderate total complexity (46.5th percentile). (d) Distribution of asymmetry across all positions. The mean asymmetry is $\langle A\rangle=-0.25$ bits (and standard deviation $1.42$ bits), suggesting a weak structural advantage for Black across Chess960 configurations. Standard chess lies at the 91.2th percentile for asymmetry.
The information cost framework reveals significant variation in decision-making burden between White and Black across Chess960 positions. We define the asymmetry as
$A=S_{B}(n)-S_{W}(n)$ (5)
which quantifies the differential complexity: positive values indicate Black faces more difficult decisions, while negative values indicate White’s burden is greater. Analysis of all 960 starting positions (Fig. 4(d)) reveals a mean asymmetry of $\langle A\rangle=-0.25$ bits (and standard deviation $\pm 0.69$ bits), suggesting a weak but systematic bias wherein White generally faces slightly more complex decision trees. This negative mean may reflect the fundamental asymmetry inherent in chess: White moves first and must immediately confront the full complexity of opening choices, while Black’s initial responses are partially constrained by White’s play. The substantial intra-position standard deviation of $1.42$ bits (from 2.51-2.51 to +1.84+1.84 bits) indicates high game-to-game variability, reflecting the sensitivity of decision complexity to specific move sequences.
II.2.2 Evolution of asymmetry during the game
Beyond the final accumulated asymmetry, it is instructive to examine how the decision complexity imbalance develops move by move during the opening phase. Figure 5 presents the complexity asymmetry $A(n)=S_{B}(n)-S_{W}(n)$ as a function of ply $n$ (the number of moves made by each player) for all 960 Chess960 starting positions.
Figure 5: Evolution of the complexity asymmetry $A(n)=S_{B}(n)-S_{W}(n)$ as a function of ply $n$ for all 960 Chess960 starting positions. Gray lines show the trajectories of individual positions (each averaged over multiple games). The blue shaded region represents the ensemble mean (blue line) $\pm 1\sigma$. The green line with markers highlights position #518 (standard chess), with its corresponding shaded band indicating $\pm 1\sigma$. Parameters: $\Delta_{0}=10$ cp, depth 20, and $n=10$ moves per player.
At $n=0$, all positions begin with $A=0$ by definition—no moves have been made, so no decisions have yet been evaluated. As play progresses, the asymmetry trajectories diverge rapidly, forming a spreading fan of possibilities. By ply $n=10$, the ensemble spans a range of approximately [2.7,+1.8][-2.7,+1.8] bits, with mean $\langle A(10)\rangle=-0.37$ bits and standard deviation $\sigma_{A}=0.74$ bits. These values are consistent with the results shown in Fig. 4, although not identical due to the different engine depths used in the two analyses. The negative mean indicates that, on average across all Chess960 positions, White faces slightly more difficult decisions than Black during the opening phase.
The roughly linear decrease of the mean asymmetry with ply indicates that, on average, White faces increasingly harder decisions as the opening progresses. The widening spread shows that positions differ not only in the magnitude of their asymmetry but also in how quickly it develops: some configurations concentrate the complexity difference in the first few moves, while others accumulate it more gradually. The trajectories confirm that asymmetry is a dynamic property—some positions remain close to zero throughout, whereas others exhibit steadily growing imbalances. The progressively widening uncertainty band reflects both true variation across positions and game-to-game fluctuations for each configuration.
Standard chess (position #518) follows a distinctive trajectory that rises above the ensemble mean, reaching $\langle A(10)\rangle=+0.65\pm 1.57$ bits at the 92.1th percentile. This positive asymmetry indicates that Black accumulates greater decision complexity than White as the opening unfolds—a pattern that emerges consistently despite the large game-to-game variance. The trajectory’s upward drift becomes apparent by approximately ply 3–4, suggesting that the asymmetry in standard chess is not an artifact of the very first moves but rather a cumulative effect of the classical piece arrangement.
From a practical standpoint, these trajectories illuminate why certain Chess960 positions may feel more or less balanced during actual play. Positions with steep asymmetry growth place one player under cognitive pressure early in the game, potentially influencing time management and error rates. The observation that standard chess exhibits above-average asymmetry growth favoring Black’s burden aligns with the classical intuition that Black must work harder to equalize, though here we quantify this imbalance in information-theoretic terms rather than positional evaluation.
II.3 Key Positions in the Chess960 Complexity Landscape
II.3.1 Extreme positions and standard chess
The information cost analysis allows us to identify outlier configurations that lie at the extremes of the asymmetry and complexity spectra (Fig. 4). It should be noted that both the engine depth and the number of games used for averaging naturally influence these specific configurations; we present them here primarily to illustrate the broad diversity of situations that arise within the chess960 landscape.
The maximum asymmetry ‘favoring’ Black complexity is position #504 (RBQNBKNR): This configuration exhibits $\langle A\rangle=+1.84\pm 1.75$ bits, with Black experiencing $\langle S_{B}\rangle=5.47\pm 1.42$ bits compared to White’s $\langle S_{W}\rangle=3.63\pm 1.02$ bits. The back-rank arrangement places the rook on the usual a and h-files, bishops on b and e files flanking the queen on c, with the king on f. This seems to create a tactical environment where Black’s responses demand more evaluation depth than White’s developmental moves (the principal variation for white appears to be the quite natural 1.d4, 2.c4, 3.Nc3).
Conversely, the position #481 (BQRBNKNR) is the most complex for White: with $\langle A\rangle=-2.51\pm 1.65$ bits, this configuration places the heaviest burden on White ($\langle S_{W}\rangle=5.96\pm 1.35$ bits versus Black’s $\langle S_{B}\rangle=3.45\pm 0.94$ bits). The bishop on the a-file, queen on b, and rook on c create immediate tactical complications for White that constrain opening choices far more severely than Black’s (the main openings seem to revolve around a4, b4, c4, Ngf3 with no clear variation emerging).
The position #89 (NNRBBKRQ) achieves near-perfect symmetry (with the depth and number of games considered here) with $\langle A\rangle=-0.002\pm 1.65$ bits and $\langle S_{W}\rangle=4.47\pm 1.22$ bits, $\langle S_{B}\rangle=4.47\pm 1.11$ bits. This configuration, with knights on a and b files, rook on c, centrally paired bishops on d and e, and queen on h, creates a balanced tactical landscape where the decision complexity faced by both players is statistically indistinguishable. Note that even when the complexity is well balanced between Black and White, the Stockfish evaluation still favors White, with the principal variation beginning with one of the moves b4, Nb3, c4, ....
The ‘most complex game’, corresponding to the maximum total information cost, is obtained for position #226 (BNRQKBNR) with $\langle S_{\mathrm{tot}}\rangle=15.00\pm 1.02$ bits (99.9th percentile), with $\langle S_{W}\rangle=7.53\pm 0.56$ bits and $\langle S_{B}\rangle=7.47\pm 0.85$ bits. Remarkably, this configuration differs from the standard starting position by only a single transposition between the bishop and the rook on the a-file, yet it produces the highest overall decision complexity in the entire Chess960 ensemble. Notably, the position also exhibits near-perfect symmetry ($\langle A\rangle=-0.06$ bits), demonstrating that high complexity need not imply imbalanced cognitive demands. The relatively low standard deviations indicate that this complexity is consistently reproduced across independent game realizations.
The ‘simplest game’ is obtained for the position #316 (NBQRKRBN) with the smallest value $\langle S_{\mathrm{tot}}\rangle=4.89\pm 2.55$ bits (0th percentile), $\langle S_{W}\rangle=2.37\pm 1.83$ bits and $\langle S_{B}\rangle=2.52\pm 1.78$ bits. The queen on c-file between knight and rook, with the king centrally placed between rooks on d and f, probably limits early tactical complications. However, the large standard deviation indicates high variability—some games from this position can be considerably more complex than others.
Standard chess (position #518, RNBQKBNR) exhibits $\langle S_{W}\rangle=4.29\pm 1.22$ bits and $\langle S_{B}\rangle=4.97\pm 1.17$ bits, yielding asymmetry $\langle A\rangle=+0.69\pm 1.69$ bits and total complexity $\langle S_{\mathrm{tot}}\rangle=9.26\pm 1.69$ bits. The positive asymmetry places standard chess at the 91th percentile, indicating that Black faces moderately harder decisions than in most Chess960 configurations. However, the uncertainty is large: zero asymmetry lies within one standard deviation, so this imbalance is not statistically significant at the single-game level. The total complexity percentile of 47th indicates that standard chess exhibits moderate overall complexity—neither exceptionally simple nor complex relative to the broader Chess960 ensemble.
The scatter plot of asymmetry versus total information cost (Fig. 4c) reveals no significant correlation between these quantities, indicating that overall game complexity is independent of the White-Black balance. Highly complex games can be either symmetric or asymmetric, and conversely, simple games show comparable variation in burden distribution. This independence suggests that Chess960’s design successfully decouples two distinct dimensions of strategic depth: the total decision-making challenge and its allocation between players.
We illustrate the large variety of both complexity and evaluation in the Fig. 6 that presents both the complexity asymmetry $A$ and the initial Stockfish evaluation $E$ as functions of the Chess960 position number. This visualization highlights the substantial diversity of both complexity and initial advantage across the full ensemble of 960 starting configurations.
Figure 6: Complexity asymmetry $A(n)=S_{B}(n)-S_{W}(n)$ for $n=10$ plies (top) and initial Stockfish evaluation $E$ (bottom) as functions of Chess960 position number. The shaded region in panel (a) represents $\pm 1\sigma$ uncertainty from multi-game averaging. Gold stars mark position #518 (standard chess). Red dashed lines indicate ensemble means ($\langle A\rangle$ and $\langle E\rangle$). The position numbering follows the standard Chess960 enumeration scheme based on back-rank piece placement. Both quantities show substantial variation across positions with no obvious structure. Calculations were done at depth 14, averaged over 5 games and $\Delta_{0}=10$cp.
The asymmetry panel (Fig. 6a) displays considerable variation across positions, with values spanning approximately [2.51,+1.84][-2.51,+1.84] bits around a slightly negative mean. The shaded uncertainty band illustrates the game-to-game variability for each position; notably, this variability is comparable in magnitude to the inter-position differences, underscoring the stochastic nature of individual game trajectories. Position #518 (standard chess) lies slightly above the mean asymmetry, consistent with its 91st percentile ranking.
The evaluation panel (Fig. 6b) shows that initial Stockfish assessments cluster tightly around the mean, with most positions evaluated within $\pm 0.1$ pawns. This narrow distribution reflects the fundamental balance built into Chess960’s design constraints: all positions satisfy the same castling prerequisites and bishop color requirements, limiting gross structural imbalances. The few outlier evaluations correspond likely to positions where piece placement creates immediate tactical asymmetries.
Importantly, the two panels reveal no obvious correlated structure: positions with extreme asymmetry $A$ do not systematically coincide with extreme evaluations $E$, and vice versa. The position numbering scheme, based on the Scharnagl encoding of back-rank arrangements, could introduce some local correlations—neighboring positions often share similar piece placements—but no clear correlations or periodic structure emerges in either quantity. This likely reflects the strong sensitivity of chess to microscopic details: a change as small as moving one piece by a single square can redirect the entire course of the game, leading to dramatically different evaluations and strategic outcomes.
From a game-design perspective, these findings illuminate the properties of the classical starting position: standard chess maintains moderate complexity (47th percentile) while exhibiting above-average asymmetry favoring Black’s decision burden (91st percentile). Position #226 (BNRQKBNR) offers an interesting alternative: it achieves the highest complexity in the ensemble while maintaining near-perfect symmetry, potentially providing both maximal strategic depth and balanced competitive conditions.
II.3.2 Looking for the most balanced position
The two fairness metrics introduced above—initial evaluation $E$ and decision asymmetry AA—capture distinct aspects of competitive balance: the former measures the expected outcome advantage, while the latter quantifies the imbalance in cognitive burden between players. A natural question is whether these metrics are correlated: do positions favoring White in evaluation also impose harder decisions on Black?
Figure 7 plots the complexity asymmetry $A=S_{B}-S_{W}$ against the initial Stockfish evaluation $E$ for all 960 starting positions. The correlation is weak ($r=0.15$), indicating that evaluation advantage and decision complexity are largely independent dimensions of positional character. Positions favoring White in evaluation do not systematically impose harder decisions on Black, nor do evaluation-balanced positions guarantee symmetric cognitive demands.
Figure 7: Complexity asymmetry $A=S_{B}-S_{W}$ versus initial Stockfish evaluation $E$ (depth 30) for all 960 Chess960 starting positions. Error bars indicate one standard deviation across 5 games per position. The weak correlation ($r=0.15$) shows that evaluation balance and decision symmetry are largely independent. The gold star marks position #518 (standard chess); the green diamond indicates position #198 (QNBRKBNR), the most balanced configuration minimizing distance to the origin in the normalized $(E,A)$ plane.
This independence motivates a search for the configuration that minimizes both imbalances simultaneously. We define the most balanced position as the one minimizing the normalized distance to the origin
$d=\sqrt{\left(\frac{E}{\sigma_{E}}\right)^{2}+\left(\frac{A}{\sigma_{A}}\right)^{2}}$ (6)
where $\sigma_{E}$ and $\sigma_{A}$ are the standard deviations of evaluation and asymmetry across the ensemble, respectively. This criterion identifies position #198 (QNBRKBNR) as the most balanced configuration, with $E=+0.03$ pawns and $A=-0.03\pm 1.68$ bits—both effectively zero within measurement uncertainty. Surprisingly, this arrangement can be obtained from the standard chess position by a single transposition between the queen and the rook on the a-file. Because the initial evaluation is very small, many openings perform similarly well, including 1.b3, 1.d4, and 1.c4.
Standard chess (position #518), by contrast, lies far from this optimum: while its evaluation is typical ($E=+0.30$ pawns), its decision asymmetry places it at the 91st percentile ($A=+0.69\pm 1.69$ bits), indicating that Black faces a greater cognitive burden than in most Chess960 configurations. The existence of configurations like position #198 that achieve near-perfect balance in both metrics suggests that tournament organizers seeking maximally fair starting positions have principled alternatives to both standard chess and random selection.
Clearly, further chess-specific analysis is required to understand why this position exhibits such balanced behavior (and why the ‘extreme’ positions discussed above have their speific properties). Relevant features may include the placement of the rooks (for example, whether both are on the queenside), the king’s position and the resulting difficulty or subtlety of castling decisions, the presence of early tactical opportunities that must be avoided, the initial geometry of the bishops—whether they immediately control long diagonals and if they operate on the same flank—and the availability of viable pawn breaks in the opening. A more detailed examination of these structural elements is needed to translate the statistical findings into concrete chess insights.
III Discussion
Our study identifies two main insights linking chess to broader ideas in complex systems: the 960 starting positions form a rugged landscape of decision complexity, and small structural changes can significantly alter the balance and difficulty of the game.
First, standard chess (position #518) occupies a statistically unremarkable location in the Chess960 complexity landscape: it shows a typical initial advantage ($E=+0.30$ pawns) and a moderate total complexity (47th percentile), while its asymmetry lies in the upper tail of the distribution (91st percentile), indicating that Black faces somewhat harder decisions in the opening. This challenges the intuitive notion that centuries of play have optimized the classical starting position for maximal strategic depth. Several hypotheses may account for this observation. The classical arrangement exhibits high visual symmetry about the board’s vertical axis, making it both aesthetically pleasing and easy to remember—a practical advantage for cultural transmission. Additionally, moderate complexity may be optimal for learning, as overly simple positions lack strategic richness while highly complex ones prove too chaotic for systematic study; the classical position may thus represent an implicit compromise between depth and accessibility. Finally, it is possible that no optimization occurred at all—the classical configuration could simply reflect historical accident, frozen by tradition once chess rules stabilized in the 15th century.
Second, we find substantial heterogeneity across the Chess960 ensemble in both complexity and fairness. The asymmetry parameter $A=S_{b}-S_{w}$ spans more than 4 bits, with some positions strongly favoring White and others Black in terms of decision-making burden. This heterogeneity has practical implications: in Freestyle chess tournaments, starting positions are typically drawn at random under the assumption that all configurations are roughly equivalent. Our results challenge this assumption—random selection does not guarantee competitive fairness, as some positions impose significantly unequal cognitive demands on the two players. This finding provides a quantitative basis for more principled position selection protocols in competitive Chess960, whether by excluding outliers in asymmetry, curating balanced subsets, or weighting draws toward fairer configurations.
IV Conclusion
In this work, we introduced an information–cost framework for quantifying decision complexity in Chess960 starting positions. Our analysis shows that the 960 configurations differ markedly in both strategic depth and competitive fairness, challenging the common assumption that all starting positions are equivalent. Standard chess, despite centuries of cultural evolution, does not occupy an exceptional location in this landscape: it exhibits a typical initial advantage and moderate total complexity, while displaying above-average asymmetry in decision difficulty. This may reflect implicit historical selection for positions that balance richness with accessibility rather than maximal complexity or perfect fairness.
More broadly, our results demonstrate how concepts from information theory and statistical physics can quantify strategic complexity in deterministic decision-making systems. The framework presented here is not specific to chess; it applies to any setting in which optimal choices can be evaluated and the difficulty of distinguishing them can be measured.
These findings also open perspectives for studying the cultural evolution of board games. The fact that the classical starting position is not structurally special raises questions about the ‘phylogenetics’ of chess variants: how the traditional configuration emerged from the broader space of possibilities, and what selection pressures—cognitive, aesthetic, pedagogical, or historical—shaped its adoption. Our approach provides quantitative tools for comparing historical games (Shatranj, Xiangqi, Shogi) and for reconstructing the underlying fitness landscape that may have influenced their evolution. It also allows one to examine whether successful games cluster in particular regions of complexity space, and whether convergent evolution toward similar complexity profiles occurred independently across cultures.
Finally, the framework can be extended by incorporating higher-order information-theoretic measures—such as mutual information between successive moves—to probe the dynamical structure of openings. Applying similar methods to other randomized-start variants or to entirely different strategic games (Go, Shogi, and modern board games) would test the generality of these ideas and enable systematic cross-game comparisons of strategic complexity.
Acknowledgements.
Acknowledgments
MB thanks the GothamChess and Blitzstream channels for providing valuable food for thought about chess.

### v2

@@
 We analyze strategic complexity across all 960 Chess960 (Fischer Random Chess) starting positions.
-Stockfish evaluations show a near-universal first-move advantage for White ($\langle E\rangle=+0.30\pm 0.14$ pawns), indicating that the advantage conferred by moving first is a robust structural feature of the game.
+Stockfish evaluations show a near-universal first-move advantage for White ($\langle E\rangle=+0.30\pm 0.14$ pawns), indicating that the initiative is a robust structural feature.
 To quantify decision difficulty, we introduce an information-based measure $S(n)$ describing the cumulative information required to identify optimal moves over the first $n$ plies.
@@
 We focus on the Grandmaster level here and set $\Delta_{0}=10$ cp.
-[…]
 Figure 4: Information cost analysis across the 960 Chess960 starting positions based on $n=10$ plies per player of optimal play (Stockfish 17.1, depth 14, discrimination threshold $\Delta_{0}=10$ cp).
@@
 Standard chess lies at the 91.2th percentile for asymmetry.
+[…]
 The information cost framework reveals significant variation in decision-making burden between White and Black across Chess960 positions. We define the asymmetry as
@@
 The substantial intra-position standard deviation of $1.42$ bits (from 2.51-2.51 to +1.84+1.84 bits) indicates high game-to-game variability, reflecting the sensitivity of decision complexity to specific move sequences.
-II.2.2 Evolution of asymmetry during the game
+II.2.2 Complexity and time to make a move
+If the quantities $\Delta$ and $S(10)$ are meaningful indicators of the intrinsic complexity of a position, one naturally expects them to correlate with the time players spend deciding on a move.
+Intuitively, positions with a small $\Delta$ (i.e., many nearly equivalent candidate moves) or a large entropy $S(10)$ (i.e., a broad distribution of plausible continuations) should require more cognitive effort, and therefore more time, to resolve [ref referee].
+To test this hypothesis, we analyze a freely available database of online games played on Lichess with move-level time stamps [28].
+For each position, we compute $\Delta$ and $S(10)$ using Stockfish at depth 15, and measure the mean time $\tau$ required by players to make a move.
+We then study the dependence of $\tau$ on $\Delta$ and $S(10)$ (Fig.
+5).
+The dataset consists of $1000$ games, corresponding to a total of 70 15070\,150 moves.
+Figure 5: (a) Mean thinking time $\tau$ as a function of $\Delta$.
+(b) Mean thinking time $\tau$ as a function of the entropy $S(10)$.
+Data are computed from $1000$ games (total of 70 15070\,150 moves) from [28], using Stockfish at depth 15.
+Since players of different strengths perceive and handle complexity differently, we account for Elo-dependent effects by adapting the reference threshold $\Delta_{0}$ as a function of the Elo rating.
+Specifically, we use
+$\Delta_{0}=\begin{cases}100,&\text{Elo}<1400,\\[4.0pt] 100-90\,\frac{\text{Elo}-1400}{1000},&1400\leq\text{Elo}\leq 2400,\\[4.0pt] 10,&\text{Elo}>2400.\end{cases}$ (6)
+This choice reflects the increasing ability of stronger players to discriminate between close alternatives and to process complex positions more efficiently.
+As expected, we observe a clear decrease of the average thinking time $\tau$ with increasing $\Delta$, and a corresponding increase of $\tau$ with $S(10)$.
+Deviations from these trends are visible for very small $\Delta$ or very large $S(10)$.
+These fluctuations can be attributed to additional factors not captured by our complexity measures, such as opening preparation: opening positions often exhibit high complexity and small $\Delta$, yet moves are played rapidly due to memorization.
+Despite these effects, the overall trends provide strong empirical support for $\Delta$ and $S(10)$ as relevant proxies for decision-making complexity in chess.
+They capture, in a quantitative manner, the intuitive notion that positions with many competing continuations or poorly differentiated evaluations are cognitively more demanding and require longer deliberation times.
+II.2.3 Evolution of asymmetry during the game
 Beyond the final accumulated asymmetry, it is instructive to examine how the decision complexity imbalance develops move by move during the opening phase.
@@
 The asymmetry panel (Fig.
-6a) displays considerable variation across positions, with values spanning approximately [2.51,+1.84][-2.51,+1.84] bits around a slightly negative mean.
+7a) displays considerable variation across positions, with values spanning approximately [2.51,+1.84][-2.51,+1.84] bits around a slightly negative mean.
 The shaded uncertainty band illustrates the game-to-game variability for each position; notably, this variability is comparable in magnitude to the inter-position differences, underscoring the stochastic nature of individual game trajectories.
@@
 The evaluation panel (Fig.
-6b) shows that initial Stockfish assessments cluster tightly around the mean, with most positions evaluated within $\pm 0.1$ pawns.
+7b) shows that initial Stockfish assessments cluster tightly around the mean, with most positions evaluated within $\pm 0.1$ pawns.
 This narrow distribution reflects the fundamental balance built into Chess960’s design constraints: all positions satisfy the same castling prerequisites and bishop color requirements, limiting gross structural imbalances.
@@
 Acknowledgments
-MB thanks the GothamChess and Blitzstream channels for providing valuable food for thought about chess.
+I thank the GothamChess and Blitzstream channels for providing valuable food for thought about chess. I also thank Giordano de Marzo for his suggestion about thinking time and complexity.

### v3

@@
 We analyze strategic complexity across all 960 Chess960 (Fischer Random Chess) starting positions.
-Stockfish evaluations show a near-universal first-move advantage for White ($\langle E\rangle=+0.30\pm 0.14$ pawns), indicating that the initiative is a robust structural feature.
-To quantify decision difficulty, we introduce an information-based measure $S(n)$ describing the cumulative information required to identify optimal moves over the first $n$ plies.
-This measure decomposes into contributions from White and Black, $S_{W}$ and $S_{B}$, yielding a total opening complexity $S_{\mathrm{tot}}=S_{W}+S_{B}$ and a decision asymmetry $A=S_{B}-S_{W}$.
-Across the ensemble, $S_{\mathrm{tot}}$ varies by a factor of three, while $A$ spans from 2.5-2.5 to +1.8+1.8 bits, showing that some openings burden White and others Black.
-The mean $\langle A\rangle=-0.25$ bits indicates a slight tendency for White to face harder opening decisions.
-Standard chess (position #518, RNBQKBNR) exhibits above-average asymmetry (91st percentile) but typical overall complexity (47th percentile).
-The most complex opening is #226 (BNRQKBNR), whereas #198 (QNBRKBNR)is the most balanced, with both evaluation and asymmetry near zero.
-These results reveal a highly heterogeneous Chess960 landscape in which small rearrangements of the back-rank pieces can significantly alter strategic depth and competitive fairness.
-Remarkably, the classical starting position—despite centuries of cultural selection—lies far from the most balanced configuration.
+Stockfish evaluations reveal a near-universal first-move advantage for White ($\langle E\rangle=+0.33\pm 0.12$ pawns), indicating that the initiative is a robust structural feature of the game.
+To quantify decision difficulty, we introduce an information-based measure $S(n)$ that captures the cumulative information required to identify optimal moves over the first $n$ plies.
+This measure decomposes into White and Black contributions, $S_{W}$ and $S_{B}$, defining a total opening complexity $S_{\mathrm{tot}}=S_{W}+S_{B}$ and a decision asymmetry $A=S_{B}-S_{W}$.
+Across the ensemble, $S_{\mathrm{tot}}$ ranges from $2.6$ to $17.2$ bits, while $A$ spans from 4.5-4.5 to +4.2+4.2 bits (mean $\langle A\rangle=-0.26$), showing that openings are nearly evenly split between those that burden White and those that burden Black, with a slight average excess complexity for White.
+Standard chess (position #518, RNBQKBNR) exhibits near-average total complexity and asymmetry, yet lies far from the configuration that jointly minimizes evaluation imbalance and decision asymmetry.
+These results reveal a highly heterogeneous Chess960 landscape in which small rearrangements of back-rank pieces can substantially alter strategic depth and competitive balance.
+The classical starting position—despite centuries of refinement—appears not as an extremum, but as one configuration among many in a broad statistical ensemble.
 I Introduction
-Chess, one of humanity’s oldest and most studied strategic games, served as a simple testing ground for theories of decision-making, artificial intelligence, and complex systems [1].
+Chess, one of humanity’s oldest and most studied strategic games, served as a simple testing ground for theories of decision-making, artificial intelligence, and complex systems Levy (1988).
 From a physics perspective, chess represents an ideal model system for studying decision-making complexity and information dynamics: it is fully deterministic, has well-defined rules, generates vast empirical datasets, and admits precise computational analysis of optimal strategies.
 The modern rules of chess crystallized in the 15th century, with the classical starting position—RNBQKBNR (Rook-Knight-Bishop-Queen-King-Bishop-Knight-Rook)—becoming the universal standard.
-This position was not derived from mathematical principles but emerged through cultural evolution and practical play over centuries [2].
+This position was not derived from mathematical principles but emerged through cultural evolution and practical play over centuries Murray (1917).
 The dominance of opening theory in modern chess has led to a paradoxical situation: at the highest levels of play, the first 15–20 moves often follow extensively analyzed variations stored in databases containing millions of games.
@@
 The tension between memorization and genuine understanding has long concerned chess theorists.
-As early as 1792, the Dutch chess enthusiast Philip Julius van Zuylen van Nijevelt proposed randomizing starting positions to restore creative play [3].
-Two centuries later, World Champion Bobby Fischer recognized the same issue as a potential threat to the game’s intellectual vitality [4].
-In 1996, Fischer proposed a variant that preserves chess’s strategic depth while eliminating memorization advantages: Fischer Random Chess, developed in collaboration with Susan Polgar [4].
-Now standardized as Chess960 or ‘Fischer random’, and more recently promoted as ‘Freestyle Chess’ [5], this variant shuffles the pieces on the back rank subject to three constraints: (i) the two bishops must occupy opposite-colored squares, (ii) the king must be positioned between the two rooks to preserve castling rights, and (iii) White and Black must share identical mirrored arrangements.
-These rules generate exactly $960$ legal starting positions, with the classical arrangement RNBQKBNR corresponding to index #518 in the standard numbering.
-The number $960$ follows directly from elementary combinatorics.
-Labeling the back-rank squares (from a to h), the bishops can be placed on any pair of light and dark squares, giving $4\times 4=16$ combinations.
-The queen may then occupy any of the remaining six squares.
-This leaves five squares to be filled by the multiset $\{R,K,R,N,N\}$.
-With 5 squares left, there are $\binom{5}{2}=10$ ways to place the two knights; the remaining three squares must then be occupied by the two rooks and the king, with the king between the rooks, which yields a unique arrangement for these three pieces.
-Altogether, this gives $16\times 6\times 10=960$ possible starting positions.
-This deterministic construction also provides a canonical indexing scheme from 0 to $959$, known as the Scharnagl enumeration algorithm [6, 7].
-In this scheme, each starting position is uniquely encoded by three integers: (i) the bishop code, selecting one of the $16$ opposite-colored bishop pairs; (ii) the queen position, selecting one of the $6$ remaining squares after the bishops are placed; and (iii) the N5N code, indexing one of the admissible arrangements of the multiset $\{R,K,R,N,N\}$ on the final five squares under the requirement that the king lies between the two rooks.
-The mapping to an index is given by $\mathrm{idn}\;=\;(\text{bishop code})\;+\;16\times(\text{queen position})\;+\;96\times(\text{N5N code})$.
-For the classical starting position RNBQKBNR, this yields $\mathrm{idn}=6\;+\;16\times 2\;+\;96\times 5=518$ which is the standard numbering for classical chess.
-This indexing provides a compact and reproducible method for scanning all $960$ starting configurations and guarantees a one-to-one correspondence between integers in $[0,959]$ and legal Chess960 starting positions.
-The variant gained official recognition when FIDE incorporated it into the Laws of Chess in 2008 and began sanctioning world championships in 2019 [4].
-More recently, freestyle chess gained more attention with the new ‘Freestyle Chess Grand Slam Tour’ [5].
+As early as 1792, the Dutch chess enthusiast Philip Julius van Zuylen van Nijevelt proposed randomizing starting positions to restore creative play van Zuylen van Nijevelt (1792).
+Two centuries later, World Champion Bobby Fischer recognized the same issue as a potential threat to the game’s intellectual vitality Polgar (2025).
+In 1996, Fischer proposed a variant that preserves chess’s strategic depth while eliminating memorization advantages: Fischer Random Chess, developed in collaboration with Susan Polgar Polgar (2025).
+Now standardized as Chess960 or ‘Fischer random’, and more recently promoted as ‘Freestyle Chess’ 13, this variant shuffles the pieces on the back rank subject to three constraints: (i) the two bishops must occupy opposite-colored squares, (ii) the king must be positioned between the two rooks to preserve castling rights, and (iii) White and Black must share identical mirrored arrangements.
+These rules generate exactly $960$ legal starting positions (see Appendix A), with the classical arrangement RNBQKBNR corresponding to index #518 in the standard numbering.
+Chess960 gained official recognition when FIDE incorporated it into the Laws of Chess in 2008 and began sanctioning world championships in 2019 Polgar (2025).
+More recently, it received even more attention with the new ‘Freestyle Chess Grand Slam Tour’ 13.
 From the standpoint of statistical physics, Chess960 offers a unique opportunity: a discrete ensemble of 960 initial conditions evolving under identical dynamical rules, enabling systematic comparison of complexity across configurations.
@@
 These questions parallel fundamental inquiries in statistical physics concerning how initial conditions influence the complexity of dynamical trajectories.
-Recent advances in computer chess, particularly the development of engines like Stockfish capable of evaluating positions with superhuman accuracy [8], enable us to address these questions rigorously.
-The availability of large databases has enabled the emergence of a quantitative ‘science of chess,’ with studies addressing innovation dynamics [9], statistical regularities in opening frequencies [10], network properties of move sequences [11, 12, 13, 14, 15], memory effects and temporal correlations [16, 17], game complexity measures [18, 19, 20], human performance and cognition [21, 22, 23], and spatial control patterns [24, 25].
+Recent advances in computer chess, particularly the development of engines like Stockfish capable of evaluating positions with superhuman accuracy Stockfish Development Team (2024), enable us to address these questions rigorously.
+The availability of large databases has enabled the emergence of a quantitative ‘science of chess,’ with studies addressing innovation dynamics Perotti et al.
+(2013), statistical regularities in opening frequencies Blasius and Tonjes (2009), network properties of move sequences Maslov (2009); Farren et al.
+(2013); Ribeiro et al.
+(2013); Barthelemy (2025b); Cerioli et al.
+(2025), memory effects and temporal correlations Schaigorodsky et al.
+(2014, 2016), game complexity measures Atashpendar et al.
+(2016); De Marzo and Servedio (2023); Chacoma and Billoni (2024), human performance and cognition Chowdhary et al.
+(2023); Chassy and Gobet (2011); Sigman et al.
+(2010), and spatial control patterns Barthelemy (2023, 2025a).
 These works demonstrate that chess provides a rich empirical system for studying decision-making under uncertainty.
@@
 II Results
-II.1 Simple measures
-II.1.1 Location probability
+In this section, we present three complementary indicators characterizing the 960 starting positions: (i) the positional probability distribution of pieces in the initial configuration, (ii) the engine-based structural evaluation of all positions, and (iii) a measure of the complexity burden borne by White and Black during the first moves of the game.
+While many other metrics could in principle be considered, our objective here is to focus on quantities directly related to structural imbalance and early-game decision difficulty.
+Additional measures—such as the branching factor—are briefly discussed in Appendix B, where we explain why they are less central to the questions addressed in this study.
+II.1 Location probability
 Figure 1: Probability distribution of pieces for the 960 initial positions across back-rank squares (from a to h).
@@
 By contrast, the king is more likely to appear near the central files, while the rooks are correspondingly more likely to be placed toward the sides.
-As discussed above, there are 960 possibilities that result from the following constraints: The two bishops must occupy opposite-colored squares (one on a light square, one on a dark square); the king must be positioned between the two rooks, preserving the ability to castle kingside and queenside.
-Finally, we note that the Black pieces mirror the White arrangement, ensuring a priori fairness.
-These 960 positions are numbered from 0 to 959 and the standard chess position (RNBQKBNR) corresponds to the number 518 in this numbering [26].
+The 960 possibilities result from the following constraints (see Appendix A): The two bishops must occupy opposite-colored squares (one on a light square, one on a dark square); the king must be positioned between the two rooks, preserving the ability to castle kingside and queenside.
+Finally, Black pieces mirror the White arrangement, ensuring a priori fairness.
+These 960 positions are numbered from 0 to 959 and the standard chess position (RNBQKBNR) corresponds to the number 518 in this numbering 9.
 The simplest question we can ask under these constraints is: what is the probability that a given piece occupies a particular file?
@@
 1 shows the probability distribution for each piece type across the eight squares, revealing the symmetry and constraints of the Chess960 rules.
-II.1.2 Initial Stockfish evaluation
-To quantify the structural asymmetry inherent in Chess960’s turn-based mechanics, we evaluated all 960 starting positions using Stockfish 17.1 [8] at depth 30.
+II.2 Initial Stockfish evaluation
+In order to quantify the intrinsic structural asymmetry of Chess960 starting positions, we perform a systematic engine evaluation of all 960 configurations.
+The goal is not to model human perception or practical playing strength, but to estimate the underlying structural advantage encoded in each initial arrangement under near-optimal continuation.
+To this end, we evaluated all 960 starting positions using Stockfish 17.1 Stockfish Development Team (2024) at fixed depth 30.
+The depth-30 score is interpreted here as an engine-level proxy for structural advantage, reflecting the evaluation after a sufficiently deep search while remaining computationally tractable.
+All computations were performed with NNUE enabled, using a single search thread (Threads=1) and a 1,GB hash table (Hash=1024).
+Each position was analyzed under a strict depth constraint without time limits, ensuring that every evaluation reached the prescribed search depth exactly.
 The resulting distribution (Fig.
-2) reveals a striking uniformity: the mean initial evaluation is $\langle E\rangle=+0.297\pm 0.136$ pawns, with 956 of 960 positions (99.6%) exhibiting positive evaluation favoring White.
-Only four position achieve exact or close to balance.
+2) reveals a striking uniformity: the mean initial evaluation is $\langle E\rangle=+0.33\pm 0.12$ pawns, with 959 of 960 positions (99.9%) exhibiting positive evaluation favoring White.
+Only one position (#774) achieves a small avantage for black (E=0.18E=-0.18).
 This near-universal first-move advantage confirms that the privilege of moving first confers a measurable strategic benefit independent of piece configuration.
-Standard chess (position #518, RNBQKBNR) exhibits an initial evaluation of +0.30+0.30 pawns, placing it in the 47th percentile.
+Standard chess (position #518, RNBQKBNR) exhibits an initial evaluation of +0.28+0.28 pawns, placing it in the 37th percentile.
 This demonstrates that the classical starting position is statistically typical within the Chess960 landscape, neither amplifying nor diminishing the structural benefit of initiative.
-Figure 2: Distribution of Stockfish initial position evaluations across all 960 Chess960 starting positions (Stockfish 17.1, depth 30).
-The distribution is centered at $\langle E\rangle=+0.30\pm 0.14$ pawns, with 99.6% of positions favoring White.
-Vertical lines indicate position #518 (standard chess, solid green), position #279 (maximum White advantage, red dashed), position #535 (almost perfect balance, blue dashed), and the ensemble mean (black dashed).
-The distribution’s relatively narrow standard deviation ($\sigma=0.14$ pawns) indicates remarkable consistency across disparate piece arrangements, suggesting that White’s first-move advantage is a robust structural feature largely independent of the specific configuration.
-The most White-favorable position (#279, NRBKNRQB) reaches +0.83+0.83 pawns—approximately 2.8 times the mean.
-This configuration places in particular the rook, the queen and a bishop on the f-h files, creating immediate developmental asymmetries that White can exploit (although the engine’s sequence of best moves starting with f4, c4, Nf3 leads to variations that are not obvious).
-Notably, position #535 (RNBKQNRB) achieves almost perfect balance (with an infinitesimal advantage for Black at 0.09-0.09 pawns) —one of only three configurations where the first-move advantage is close to null.
-The rarity of balanced or Black-favoring configurations (0.4% of positions) underscores that neutralizing White’s first-move advantage requires a precise alignment of spatial and tactical factors.
-Note that these results depend on the engine and search depth.
-While specific outliers may shift, the overall distribution and its statistics remain robust.
+Figure 2: Distribution of Stockfish initial evaluations across all 960 Chess960 starting positions (Stockfish 17.1, depth 30, Threads=1).
+The distribution is centered at $\langle E\rangle=+0.33\pm 0.12$ pawns, with $99.9\%$ of positions favoring White.
+Vertical lines indicate position #518 (standard chess, solid green).
+At this depth, position #333 exhibits the maximum White advantage (red dashed), while position #774 is the closest to perfect evaluation balance (blue dashed).
+The ensemble mean is shown as a black dashed line.
+The distribution’s relatively narrow standard deviation ($\sigma=0.12$ pawns) indicates remarkable consistency across disparate piece arrangements, suggesting that White’s first-move advantage is a robust structural feature largely independent of the specific configuration.
+The most White-favorable position (#333, NRQBKRBN) reaches the surprisingly large value of +0.81+0.81 pawns—approximately 2.5 times the mean.
+Inspection of the principal variation reveals that White can rapidly mobilize its pieces with the plan Nb3, f4, and Bc5, seizing central and kingside space while targeting the weakened dark squares around Black’s king.
+The structural origin of this asymmetry might lie in the initial placement itself: White’s queen on c1 and bishop on d1 are already aligned along natural attacking diagonals, while the knight on a1 finds an efficient route to b3 where it supports both the center and the c5 outpost.
+More analysis of the variations are however necessary to inspect in more details this asymmetry.
+Another interesting case is the only position #774 (QRBKNBRN) that gives an advantage to black, although very small (0.18-0.18).
+The rarity of balanced configurations underscores that neutralizing White’s first-move advantage requires a precise alignment of spatial and tactical factors.
 From a game-theoretic perspective, the universality of White’s advantage across Chess960 positions suggests that the first-move privilege is not an artifact of centuries of opening theory refinement in standard chess, but rather an intrinsic property of the game’s mechanics.
@@
 This finding has implications for tournament fairness: while Chess960 successfully neutralizes preparation advantages, it preserves the need for color-balanced pairings or double-game formats to ensure equity between opponents.
-II.1.3 Branching factor and position space growth analysis
-To quantify the complexity evolution across different Chess960 starting positions, we analyze two key metrics: the branching factor $b(n)$ and the position space size $N(n)$ (i.e.
-the number of different positions) as functions of ply depth $n$.
-This analysis characterizes the raw combinatorial complexity of the game tree, independent of strategic considerations or optimal play.
-It quantifies the fundamental decision-making burden at each stage of the game: how many distinct game states a player must consider when planning ahead.
-For each position, we compute $N(n)$ defined as the number of unique legal positions reachable at ply depth $n$.
-We then compute the branching factor $b(n)$ defined as the ratio of consecutive position counts
-$b(n)=\frac{N(n+1)}{N(n)}$ (1)
-This ratio quantifies the average number of accessible successor states per position at depth $n$.
-Figure 3 presents a comparative analysis between position 518 (standard chess) and the ensemble average across all sampled positions.
-Panel (a) shows the evolution of the branching factor $b(n)$ versus ply depth.
-The average branching factor across sampled positions (blue curve with error bars representing one standard deviation) exhibits typical values of order $20$ consistent with previous estimates [24].
-Note that branching factors around $b\sim 35$, often cited in computer chess literature, correspond to middlegame positions where piece development leads to increased move options [24].
-Position #518 (green curve) closely tracks this average, with deviations remaining within one standard deviation throughout the analyzed range.
-Panel (b) displays the growth of unique positions $N(n)$ on a semi-logarithmic scale.
-The approximately linear behavior confirms exponential growth, $N(n)\sim\exp(\lambda n)$.
-The extremely small dispersion around the average (smaller than the symbol size) indicates that all starting positions generate state spaces of essentially identical size.
-Note that the effective growth rate $\lambda$ is notably smaller than what would be predicted from the branching factor alone ($\lambda<\ln\langle b\rangle$).
-This reduction arises because different move sequences can converge to the same board configuration—a phenomenon known as transposition in chess terminology—causing the position space to grow more slowly than the space of legal move sequences.
-Figure 3: (a) Average branching factor $\langle b\rangle$ versus ply depth $n$ across all 960 starting configurations (blue), compared with standard chess (position #518, green).
-Error bars indicate one standard deviation.
-The dashed line marks the commonly cited middlegame value $b\approx 35$.
-(b) Number of unique positions $N(n)$ versus ply depth on a semi-logarithmic scale, showing exponential growth $N(n)\sim b_{\mathrm{eff}}^{n}$ with effective branching factor $b_{\mathrm{eff}}\approx 14.41$, significantly smaller than $\langle b\rangle$ due to transpositions.
-From the perspective of branching factors and the number of unique positions, our results show no significant dependence on the initial arrangement; in this sense, all 960 starting positions are equivalent.
-II.2 Characterizing the opening complexity
-II.2.1 Asymmetry in Decision Complexity
+II.3 Effect of depth: instability of extremes
+The outliers identified in the previous section were obtained at depth $30$.
+A natural question is whether these extreme evaluations remain stable as the search depth varies.
+In Fig.
+3, we plot the three positions with the largest evaluations and the three with the smallest evaluations (as previously computed with Stockfish 17.1, Threads=1, NNUE enabled, identical hardware and hash settings) as a function of search depth.
+Figure 3: (Top) Evaluation of the three highest- and three lowest-ranked initial positions as a function of search depth.
+(Bottom) Evaluation of position #333 as a function of depth.
+All evaluations are expressed in pawns from White’s perspective.
+The result is clear: extreme positions are not depth-stable.
+While aggregate statistics (mean, variance, overall distribution shape) remain robust, individual outliers fluctuate substantially, as illustrated in Fig.
+3(top).
+At shallow depth the instability is striking.
+At depth 10, the top configuration (position 702) reaches an evaluation of +5.89+5.89 pawns, followed by +3.41+3.41 and +2.19+2.19.
+These extreme values completely disappear at greater depth.
+Already at depth 15, the top evaluations drop to $\approx+1.16$, and by depth 30 the maximum is +0.81+0.81.
+At depth 40, the leading positions cluster tightly around +0.79+0.79.
+Thus, the apparent large advantages observed at depth 10 are largely artifacts of shallow search.
+For bottom positions variations are much smaller.
+At depth 10, the lowest evaluations are only mildly negative (0.21-0.21, 0.16-0.16, 0.10-0.10), but the identities of these configurations change repeatedly with depth.
+For instance, position 774 appears among the bottom three at depths 30 and 40, while other initially weak positions disappear from the extreme tail.
+Moreover, several configurations that are slightly negative at shallow depth become approximately neutral at intermediate depths, before stabilizing again.
+The turnover in the ranking is equally informative.
+None of the top three positions at depth 10 remains in the top three at depth 20 or beyond.
+Conversely, position #333—identified as the strongest configuration at depth 30 (+0.81+0.81)—does not appear among the leaders at shallow depth.
+Even between depths 30 and 40, although the numerical evaluations become closer, the composition of the top group still changes.
+To illustrate this sensitivity, we focus specifically on position #333 and display in Fig.
+3 its evaluation as a function of search depth.
+The figure confirms that the evaluation is not stable with depth and does not exhibit clear convergence within the explored range.
+While the advantage peaks around depths 25–30, it subsequently decreases, reaching approximately +0.5+0.5 pawns at depth 40.
+This behavior indicates that extreme evaluations are particularly sensitive to deeper tactical or strategic resources revealed by extended search.
+Positions that appear highly favorable at shallow depth may rely on lines that are later mitigated by longer forcing sequences or defensive resources uncovered at greater depth.
+More generally, this raises two questions: (i) do some positions converge rapidly and remain stable across depths, while others exhibit strong depth sensitivity?
+and (ii) which structural features control this stability?
+Addressing these issues would require a systematic study and the introduction of a quantitative measure of evaluation stability, allowing positions to be classified according to their sensitivity to search-horizon effects.
+In summary, while global evaluation statistics are robust, the ranking of extreme Chess960 starting positions is not invariant with respect to search depth.
+Claims about the “most advantageous” or “most disadvantageous” configurations should therefore be supported by an explicit depth-stability analysis.
+Search depth should be interpreted as defining the resolution at which the decision landscape is probed.
+II.4 Characterizing the opening complexity
 The measures discussed so far, however, do not quantify how difficult it is to play a given position as White or Black, nor do they assess whether one starting configuration is intrinsically more complex than another.
@@
 Consider a position where a chess engine evaluates (in centipawns) the best move with score $E_{1}$ and the second-best move with score $E_{2}$.
-The difference $\Delta=E_{1}-E_{2}$ quantifies how much ‘better’ the optimal move is [24, 25].
-For a player with discrimination ability $\Delta_{0}$ (the minimum evaluation difference they can reliably perceive), we model the probability of choosing the optimal move (over the second best move) using a Boltzmann-like factor (also called the softmax function)
+The difference $\Delta=E_{1}-E_{2}$ quantifies how much ‘better’ the optimal move is Barthelemy (2023, 2025a).
+For a player with discrimination ability $\Delta_{0}$ (the minimum evaluation difference they can reliably perceive), we model the probability $P(\text{optimal})$ of choosing the optimal move (over the second best move) using a Boltzmann-like factor (also called the softmax function)
 $P(\text{optimal})=\frac{1}{1+e^{-\Delta/\Delta_{0}}}$ (1)
@@
   • If $\Delta\gg\Delta_{0}$, the best move is effectively forced: the information cost collapses to zero, $S(\Delta)\approx 0$ bit.
-For a trajectory of $n$ plies, the cumulative information required to navigate the resulting branch of the search tree is
+For a trajectory of $n$ plies, the cumulative information required to navigate the corresponding branch of the search tree is
 $S(n)=\sum_{i=1}^{n}S(\Delta_{i})=\sum_{i=1}^{n}\log_{2}\left(1+e^{-\Delta_{i}/\Delta_{0}}\right),$ (3)
-with $\Delta_{i}$ the evaluation gap at ply $i$.
-The average information per move, $S(n)/n$, reflects the typical decision difficulty encountered along the trajectory.
-$S(n)$ provides a principled, information-theoretic measure of how demanding a given sequence of moves (such as an opening for example) is to navigate under optimal play.
-The quantity $S(n)$ has several natural properties: it is monotonically increasing with $n$ (information accumulates along the line of play); forced positions with large $\Delta_{i}$ contribute negligibly; subtle positions with small $\Delta_{i}$ dominate; and the overall magnitude depends on playing strength through $\Delta_{0}$.
-Typical values are $\Delta_{0}\approx 10$ cp for Grandmaster level, $\Delta_{0}\approx 50$ cp for Expert level, and $\Delta_{0}\approx 100$ cp for intermediate level [25].
-We focus on the Grandmaster level here and set $\Delta_{0}=10$ cp.
-Figure 4: Information cost analysis across the 960 Chess960 starting positions based on $n=10$ plies per player of optimal play (Stockfish 17.1, depth 14, discrimination threshold $\Delta_{0}=10$ cp).
-Values represent averages over 5 games per position.
-(a) Distribution of White’s cumulative cost $S_{W}(10)$ showing the decision complexity faced by White across different starting configurations.
-The green vertical line marks position #518 (standard chess) at $\langle S_{W}\rangle=4.29$ bits, near the ensemble mean.
+where $\Delta_{i}$ denotes the evaluation gap at ply $i$.
+The average information per move, $S(n)/n$, characterizes the typical decision difficulty encountered along the trajectory.
+Thus, $S(n)$ provides an information-theoretic measure of how demanding it is to follow a given sequence of moves (for example, an opening line) under optimal play.
+The quantity $S(n)$ has several natural properties.
+It is monotonically increasing with $n$, reflecting the cumulative nature of information along the line of play.
+Positions with large $\Delta_{i}$ (quasi-forced moves) contribute negligibly, whereas positions with small $\Delta_{i}$ dominate the sum.
+The overall scale of $S(n)$ depends on the discrimination parameter $\Delta_{0}$, which encodes the sensitivity to small evaluation differences.
+In Appendix C, we further show that $S(10)$ correlates with the mean thinking time $\tau$ of human players: $\tau$ increases with $S(10)$, supporting its interpretation as a meaningful proxy for decision-making complexity.
+Importantly, $S(n)$ is not intended to model human computational limits.
+Rather, it quantifies the intrinsic difficulty of identifying the optimal path among competing continuations (here in the early game).
+In this sense, $S(n)$ is a structural property of the position: it measures how sharply the optimal move is separated from alternatives along the principal variation.
+It is therefore independent of player strength and instead characterizes the geometry of the decision landscape itself.
+Estimating a precise value for $\Delta_{0}$ is non-trivial, and any choice is necessarily somewhat arbitrary.
+In Appendix C, we show that the percentile ranking of a representative position (the standard position #518) varies only weakly with $\Delta_{0}$, indicating robustness with respect to this parameter.
+We also provide empirical evidence that $\Delta_{0}\approx 10$ cp corresponds to a natural discrimination scale at expert (Grandmaster) level.
+In what follows, we therefore focus on this high-skill regime and set $\Delta_{0}=10$ cp.
+Figure 4: Information cost analysis across the 960 Chess960 starting positions, computed for $n=10$ plies per player under optimal play (Stockfish 17.1, depth 15, discrimination threshold $\Delta_{0}=10$ cp).
+(a) Distribution of White’s cumulative cost $S_{W}(10)$, quantifying the decision complexity faced by White across starting configurations.
+The green vertical line marks position #518 (standard chess), with $\langle S_{W}\rangle=5.42$ bits.
 (b) Distribution of Black’s cumulative cost $S_{B}(10)$.
-Position #518 exhibits $\langle S_{B}\rangle=4.97$ bits, slightly above the ensemble mean.
+Position #518 yields $\langle S_{B}\rangle=5.79$ bits, slightly above the ensemble mean.
 (c) Asymmetry $A=S_{B}-S_{W}$ versus total information cost $S_{\mathrm{tot}}=S_{W}+S_{B}$.
-Negative asymmetry indicates White faces more complex decisions; positive values indicate Black faces greater complexity.
-Standard chess (position #518, indicated by a gold star) shows positive asymmetry ($\langle A\rangle=+0.69$ bits) and moderate total complexity (46.5th percentile).
+Negative asymmetry indicates greater decision complexity for White; positive values indicate greater complexity for Black.
+The average asymmetry is slightly negative, $\langle A\rangle=-0.26$ bits (standard deviation $1.36$), while the average total cost is $\langle S_{\mathrm{tot}}\rangle=9.69$ bits (standard deviation $2.4$).
+Standard chess (position #518, gold star) exhibits positive asymmetry ($A_{518}=+0.37$ bits) and a relatively high total cost ($11.2$ bits, 72.5th percentile).
 (d) Distribution of asymmetry across all positions.
-The mean asymmetry is $\langle A\rangle=-0.25$ bits (and standard deviation $1.42$ bits), suggesting a weak structural advantage for Black across Chess960 configurations.
-Standard chess lies at the 91.2th percentile for asymmetry.
+The mean asymmetry is $\langle A\rangle=-0.26$ bits, indicating a slight average structural complexity advantage for Black across Chess960.
+Standard chess lies at the 69th percentile of the asymmetry distribution, placing it close to the ensemble average.
 […]
-The information cost framework reveals significant variation in decision-making burden between White and Black across Chess960 positions. We define the asymmetry as
+The information cost framework reveals important variations in decision-making burden between White and Black across Chess960 positions. We define the asymmetry as
 $A=S_{B}(n)-S_{W}(n)$ (4)
@@
 Analysis of all 960 starting positions (Fig.
-4(d)) reveals a mean asymmetry of $\langle A\rangle=-0.25$ bits (and standard deviation $\pm 0.69$ bits), suggesting a weak but systematic bias wherein White generally faces slightly more complex decision trees.
+4(d)) reveals a mean asymmetry of $\langle A\rangle=-0.26$ bits (and standard deviation $\pm 1.36$ bits), suggesting a weak but systematic bias wherein White generally faces slightly more complex decision trees.
 This negative mean may reflect the fundamental asymmetry inherent in chess: White moves first and must immediately confront the full complexity of opening choices, while Black’s initial responses are partially constrained by White’s play.
-The substantial intra-position standard deviation of $1.42$ bits (from 2.51-2.51 to +1.84+1.84 bits) indicates high game-to-game variability, reflecting the sensitivity of decision complexity to specific move sequences.
-II.2.2 Complexity and time to make a move
-If the quantities $\Delta$ and $S(10)$ are meaningful indicators of the intrinsic complexity of a position, one naturally expects them to correlate with the time players spend deciding on a move.
-Intuitively, positions with a small $\Delta$ (i.e., many nearly equivalent candidate moves) or a large entropy $S(10)$ (i.e., a broad distribution of plausible continuations) should require more cognitive effort, and therefore more time, to resolve [ref referee].
-To test this hypothesis, we analyze a freely available database of online games played on Lichess with move-level time stamps [28].
-For each position, we compute $\Delta$ and $S(10)$ using Stockfish at depth 15, and measure the mean time $\tau$ required by players to make a move.
-We then study the dependence of $\tau$ on $\Delta$ and $S(10)$ (Fig.
-5).
-The dataset consists of $1000$ games, corresponding to a total of 70 15070\,150 moves.
-Figure 5: (a) Mean thinking time $\tau$ as a function of $\Delta$.
-(b) Mean thinking time $\tau$ as a function of the entropy $S(10)$.
-Data are computed from $1000$ games (total of 70 15070\,150 moves) from [28], using Stockfish at depth 15.
-Since players of different strengths perceive and handle complexity differently, we account for Elo-dependent effects by adapting the reference threshold $\Delta_{0}$ as a function of the Elo rating.
-Specifically, we use
-$\Delta_{0}=\begin{cases}100,&\text{Elo}<1400,\\[4.0pt] 100-90\,\frac{\text{Elo}-1400}{1000},&1400\leq\text{Elo}\leq 2400,\\[4.0pt] 10,&\text{Elo}>2400.\end{cases}$ (6)
-This choice reflects the increasing ability of stronger players to discriminate between close alternatives and to process complex positions more efficiently.
-As expected, we observe a clear decrease of the average thinking time $\tau$ with increasing $\Delta$, and a corresponding increase of $\tau$ with $S(10)$.
-Deviations from these trends are visible for very small $\Delta$ or very large $S(10)$.
-These fluctuations can be attributed to additional factors not captured by our complexity measures, such as opening preparation: opening positions often exhibit high complexity and small $\Delta$, yet moves are played rapidly due to memorization.
-Despite these effects, the overall trends provide strong empirical support for $\Delta$ and $S(10)$ as relevant proxies for decision-making complexity in chess.
-They capture, in a quantitative manner, the intuitive notion that positions with many competing continuations or poorly differentiated evaluations are cognitively more demanding and require longer deliberation times.
-II.2.3 Evolution of asymmetry during the game
-Beyond the final accumulated asymmetry, it is instructive to examine how the decision complexity imbalance develops move by move during the opening phase.
-Figure 6 presents the complexity asymmetry $A(n)=S_{B}(n)-S_{W}(n)$ as a function of ply $n$ (the number of moves made by each player) for all 960 Chess960 starting positions.
-Figure 6: Evolution of the complexity asymmetry $A(n)=S_{B}(n)-S_{W}(n)$ as a function of ply $n$ for all 960 Chess960 starting positions.
-Gray lines show the trajectories of individual positions (each averaged over multiple games).
-The blue shaded region represents the ensemble mean (blue line) $\pm 1\sigma$.
-The green line with markers highlights position #518 (standard chess), with its corresponding shaded band indicating $\pm 1\sigma$.
-Parameters: $\Delta_{0}=10$ cp, depth 20, and $n=10$ moves per player.
-At $n=0$, all positions begin with $A=0$ by definition—no moves have been made, so no decisions have yet been evaluated.
-As play progresses, the asymmetry trajectories diverge rapidly, forming a spreading fan of possibilities.
-By ply $n=10$, the ensemble spans a range of approximately [2.7,+1.8][-2.7,+1.8] bits, with mean $\langle A(10)\rangle=-0.37$ bits and standard deviation $\sigma_{A}=0.74$ bits.
-These values are consistent with the results shown in Fig.
-4, although not identical due to the different engine depths used in the two analyses.
-The negative mean indicates that, on average across all Chess960 positions, White faces slightly more difficult decisions than Black during the opening phase.
-The roughly linear decrease of the mean asymmetry with ply indicates that, on average, White faces increasingly harder decisions as the opening progresses.
-The widening spread shows that positions differ not only in the magnitude of their asymmetry but also in how quickly it develops: some configurations concentrate the complexity difference in the first few moves, while others accumulate it more gradually.
-The trajectories confirm that asymmetry is a dynamic property—some positions remain close to zero throughout, whereas others exhibit steadily growing imbalances.
-The progressively widening uncertainty band reflects both true variation across positions and game-to-game fluctuations for each configuration.
-Standard chess (position #518) follows a distinctive trajectory that rises above the ensemble mean, reaching $\langle A(10)\rangle=+0.65\pm 1.57$ bits at the 92.1th percentile.
-This positive asymmetry indicates that Black accumulates greater decision complexity than White as the opening unfolds—a pattern that emerges consistently despite the large game-to-game variance.
-The trajectory’s upward drift becomes apparent by approximately ply 3–4, suggesting that the asymmetry in standard chess is not an artifact of the very first moves but rather a cumulative effect of the classical piece arrangement.
-From a practical standpoint, these trajectories illuminate why certain Chess960 positions may feel more or less balanced during actual play.
-Positions with steep asymmetry growth place one player under cognitive pressure early in the game, potentially influencing time management and error rates.
-The observation that standard chess exhibits above-average asymmetry growth favoring Black’s burden aligns with the classical intuition that Black must work harder to equalize, though here we quantify this imbalance in information-theoretic terms rather than positional evaluation.
-II.3 Key Positions in the Chess960 Complexity Landscape
-II.3.1 Extreme positions and standard chess
-The information cost analysis allows us to identify outlier configurations that lie at the extremes of the asymmetry and complexity spectra (Fig.
-4).
-It should be noted that both the engine depth and the number of games used for averaging naturally influence these specific configurations; we present them here primarily to illustrate the broad diversity of situations that arise within the chess960 landscape.
-The maximum asymmetry ‘favoring’ Black complexity is position #504 (RBQNBKNR): This configuration exhibits $\langle A\rangle=+1.84\pm 1.75$ bits, with Black experiencing $\langle S_{B}\rangle=5.47\pm 1.42$ bits compared to White’s $\langle S_{W}\rangle=3.63\pm 1.02$ bits.
-The back-rank arrangement places the rook on the usual a and h-files, bishops on b and e files flanking the queen on c, with the king on f.
-This seems to create a tactical environment where Black’s responses demand more evaluation depth than White’s developmental moves (the principal variation for white appears to be the quite natural 1.d4, 2.c4, 3.Nc3).
-Conversely, the position #481 (BQRBNKNR) is the most complex for White: with $\langle A\rangle=-2.51\pm 1.65$ bits, this configuration places the heaviest burden on White ($\langle S_{W}\rangle=5.96\pm 1.35$ bits versus Black’s $\langle S_{B}\rangle=3.45\pm 0.94$ bits).
-The bishop on the a-file, queen on b, and rook on c create immediate tactical complications for White that constrain opening choices far more severely than Black’s (the main openings seem to revolve around a4, b4, c4, Ngf3 with no clear variation emerging).
-The position #89 (NNRBBKRQ) achieves near-perfect symmetry (with the depth and number of games considered here) with $\langle A\rangle=-0.002\pm 1.65$ bits and $\langle S_{W}\rangle=4.47\pm 1.22$ bits, $\langle S_{B}\rangle=4.47\pm 1.11$ bits.
-This configuration, with knights on a and b files, rook on c, centrally paired bishops on d and e, and queen on h, creates a balanced tactical landscape where the decision complexity faced by both players is statistically indistinguishable.
-Note that even when the complexity is well balanced between Black and White, the Stockfish evaluation still favors White, with the principal variation beginning with one of the moves b4, Nb3, c4, ....
-The ‘most complex game’, corresponding to the maximum total information cost, is obtained for position #226 (BNRQKBNR) with $\langle S_{\mathrm{tot}}\rangle=15.00\pm 1.02$ bits (99.9th percentile), with $\langle S_{W}\rangle=7.53\pm 0.56$ bits and $\langle S_{B}\rangle=7.47\pm 0.85$ bits.
-Remarkably, this configuration differs from the standard starting position by only a single transposition between the bishop and the rook on the a-file, yet it produces the highest overall decision complexity in the entire Chess960 ensemble.
-Notably, the position also exhibits near-perfect symmetry ($\langle A\rangle=-0.06$ bits), demonstrating that high complexity need not imply imbalanced cognitive demands.
-The relatively low standard deviations indicate that this complexity is consistently reproduced across independent game realizations.
-The ‘simplest game’ is obtained for the position #316 (NBQRKRBN) with the smallest value $\langle S_{\mathrm{tot}}\rangle=4.89\pm 2.55$ bits (0th percentile), $\langle S_{W}\rangle=2.37\pm 1.83$ bits and $\langle S_{B}\rangle=2.52\pm 1.78$ bits.
-The queen on c-file between knight and rook, with the king centrally placed between rooks on d and f, probably limits early tactical complications.
-However, the large standard deviation indicates high variability—some games from this position can be considerably more complex than others.
-Standard chess (position #518, RNBQKBNR) exhibits $\langle S_{W}\rangle=4.29\pm 1.22$ bits and $\langle S_{B}\rangle=4.97\pm 1.17$ bits, yielding asymmetry $\langle A\rangle=+0.69\pm 1.69$ bits and total complexity $\langle S_{\mathrm{tot}}\rangle=9.26\pm 1.69$ bits.
-The positive asymmetry places standard chess at the 91th percentile, indicating that Black faces moderately harder decisions than in most Chess960 configurations.
-However, the uncertainty is large: zero asymmetry lies within one standard deviation, so this imbalance is not statistically significant at the single-game level.
-The total complexity percentile of 47th indicates that standard chess exhibits moderate overall complexity—neither exceptionally simple nor complex relative to the broader Chess960 ensemble.
+The substantial variation from 4.4-4.4 to +4.2+4.2 bits) indicates high game-to-game variability, reflecting the sensitivity of decision complexity to openings for specific initial conditions.
+Standard chess (position #518, RNBQKBNR) exhibits $\langle S_{W}\rangle=5.42$ bits and $\langle S_{B}\rangle=5.79$ bits, yielding asymmetry $\langle A\rangle=+0.36$ bits and total complexity $\langle S_{\mathrm{tot}}\rangle=11.2$ bits.
+The positive asymmetry places standard chess at the 69th percentile, indicating that Black faces moderately harder decisions than in most Chess960 configurations.
+The total complexity percentile of 72.5th indicates that standard chess exhibits above average overall complexity, but can be considered as neither exceptionally simple nor complex relative to the broader Chess960 ensemble.
 The scatter plot of asymmetry versus total information cost (Fig.
@@
 This independence suggests that Chess960’s design successfully decouples two distinct dimensions of strategic depth: the total decision-making challenge and its allocation between players.
-We illustrate the large variety of both complexity and evaluation in the Fig.
-7 that presents both the complexity asymmetry $A$ and the initial Stockfish evaluation $E$ as functions of the Chess960 position number.
-This visualization highlights the substantial diversity of both complexity and initial advantage across the full ensemble of 960 starting configurations.
-Figure 7: Complexity asymmetry $A(n)=S_{B}(n)-S_{W}(n)$ for $n=10$ plies (top) and initial Stockfish evaluation $E$ (bottom) as functions of Chess960 position number.
-The shaded region in panel (a) represents $\pm 1\sigma$ uncertainty from multi-game averaging.
-Gold stars mark position #518 (standard chess).
-Red dashed lines indicate ensemble means ($\langle A\rangle$ and $\langle E\rangle$).
-The position numbering follows the standard Chess960 enumeration scheme based on back-rank piece placement.
-Both quantities show substantial variation across positions with no obvious structure.
-Calculations were done at depth 14, averaged over 5 games and $\Delta_{0}=10$cp.
-The asymmetry panel (Fig.
-7a) displays considerable variation across positions, with values spanning approximately [2.51,+1.84][-2.51,+1.84] bits around a slightly negative mean.
-The shaded uncertainty band illustrates the game-to-game variability for each position; notably, this variability is comparable in magnitude to the inter-position differences, underscoring the stochastic nature of individual game trajectories.
-Position #518 (standard chess) lies slightly above the mean asymmetry, consistent with its 91st percentile ranking.
-The evaluation panel (Fig.
-7b) shows that initial Stockfish assessments cluster tightly around the mean, with most positions evaluated within $\pm 0.1$ pawns.
-This narrow distribution reflects the fundamental balance built into Chess960’s design constraints: all positions satisfy the same castling prerequisites and bishop color requirements, limiting gross structural imbalances.
-The few outlier evaluations correspond likely to positions where piece placement creates immediate tactical asymmetries.
-Importantly, the two panels reveal no obvious correlated structure: positions with extreme asymmetry $A$ do not systematically coincide with extreme evaluations $E$, and vice versa.
-The position numbering scheme, based on the Scharnagl encoding of back-rank arrangements, could introduce some local correlations—neighboring positions often share similar piece placements—but no clear correlations or periodic structure emerges in either quantity.
-This likely reflects the strong sensitivity of chess to microscopic details: a change as small as moving one piece by a single square can redirect the entire course of the game, leading to dramatically different evaluations and strategic outcomes.
-From a game-design perspective, these findings illuminate the properties of the classical starting position: standard chess maintains moderate complexity (47th percentile) while exhibiting above-average asymmetry favoring Black’s decision burden (91st percentile).
-Position #226 (BNRQKBNR) offers an interesting alternative: it achieves the highest complexity in the ensemble while maintaining near-perfect symmetry, potentially providing both maximal strategic depth and balanced competitive conditions.
-II.3.2 Looking for the most balanced position
-The two fairness metrics introduced above—initial evaluation $E$ and decision asymmetry AA—capture distinct aspects of competitive balance: the former measures the expected outcome advantage, while the latter quantifies the imbalance in cognitive burden between players.
-A natural question is whether these metrics are correlated: do positions favoring White in evaluation also impose harder decisions on Black?
-Figure 8 plots the complexity asymmetry $A=S_{B}-S_{W}$ against the initial Stockfish evaluation $E$ for all 960 starting positions.
-The correlation is weak ($r=0.15$), indicating that evaluation advantage and decision complexity are largely independent dimensions of positional character.
-Positions favoring White in evaluation do not systematically impose harder decisions on Black, nor do evaluation-balanced positions guarantee symmetric cognitive demands.
-Figure 8: Complexity asymmetry $A=S_{B}-S_{W}$ versus initial Stockfish evaluation $E$ (depth 30) for all 960 Chess960 starting positions.
-Error bars indicate one standard deviation across 5 games per position.
-The weak correlation ($r=0.15$) shows that evaluation balance and decision symmetry are largely independent.
-The gold star marks position #518 (standard chess); the green diamond indicates position #198 (QNBRKBNR), the most balanced configuration minimizing distance to the origin in the normalized $(E,A)$ plane.
-This independence motivates a search for the configuration that minimizes both imbalances simultaneously.
-We define the most balanced position as the one minimizing the normalized distance to the origin
-$d=\sqrt{\left(\frac{E}{\sigma_{E}}\right)^{2}+\left(\frac{A}{\sigma_{A}}\right)^{2}}$ (7)
-where $\sigma_{E}$ and $\sigma_{A}$ are the standard deviations of evaluation and asymmetry across the ensemble, respectively.
-This criterion identifies position #198 (QNBRKBNR) as the most balanced configuration, with $E=+0.03$ pawns and $A=-0.03\pm 1.68$ bits—both effectively zero within measurement uncertainty.
-Surprisingly, this arrangement can be obtained from the standard chess position by a single transposition between the queen and the rook on the a-file.
-Because the initial evaluation is very small, many openings perform similarly well, including 1.b3, 1.d4, and 1.c4.
-Standard chess (position #518), by contrast, lies far from this optimum: while its evaluation is typical ($E=+0.30$ pawns), its decision asymmetry places it at the 91st percentile ($A=+0.69\pm 1.69$ bits), indicating that Black faces a greater cognitive burden than in most Chess960 configurations.
-The existence of configurations like position #198 that achieve near-perfect balance in both metrics suggests that tournament organizers seeking maximally fair starting positions have principled alternatives to both standard chess and random selection.
-Clearly, further chess-specific analysis is required to understand why this position exhibits such balanced behavior (and why the ‘extreme’ positions discussed above have their speific properties).
-Relevant features may include the placement of the rooks (for example, whether both are on the queenside), the king’s position and the resulting difficulty or subtlety of castling decisions, the presence of early tactical opportunities that must be avoided, the initial geometry of the bishops—whether they immediately control long diagonals and if they operate on the same flank—and the availability of viable pawn breaks in the opening.
-A more detailed examination of these structural elements is needed to translate the statistical findings into concrete chess insights.
-III Discussion
-Our study identifies two main insights linking chess to broader ideas in complex systems: the 960 starting positions form a rugged landscape of decision complexity, and small structural changes can significantly alter the balance and difficulty of the game.
-First, standard chess (position #518) occupies a statistically unremarkable location in the Chess960 complexity landscape: it shows a typical initial advantage ($E=+0.30$ pawns) and a moderate total complexity (47th percentile), while its asymmetry lies in the upper tail of the distribution (91st percentile), indicating that Black faces somewhat harder decisions in the opening.
-This challenges the intuitive notion that centuries of play have optimized the classical starting position for maximal strategic depth.
-Several hypotheses may account for this observation.
-The classical arrangement exhibits high visual symmetry about the board’s vertical axis, making it both aesthetically pleasing and easy to remember—a practical advantage for cultural transmission.
-Additionally, moderate complexity may be optimal for learning, as overly simple positions lack strategic richness while highly complex ones prove too chaotic for systematic study; the classical position may thus represent an implicit compromise between depth and accessibility.
-Finally, it is possible that no optimization occurred at all—the classical configuration could simply reflect historical accident, frozen by tradition once chess rules stabilized in the 15th century.
-Second, we find substantial heterogeneity across the Chess960 ensemble in both complexity and fairness.
-The asymmetry parameter $A=S_{b}-S_{w}$ spans more than 4 bits, with some positions strongly favoring White and others Black in terms of decision-making burden.
-This heterogeneity has practical implications: in Freestyle chess tournaments, starting positions are typically drawn at random under the assumption that all configurations are roughly equivalent.
-Our results challenge this assumption—random selection does not guarantee competitive fairness, as some positions impose significantly unequal cognitive demands on the two players.
-This finding provides a quantitative basis for more principled position selection protocols in competitive Chess960, whether by excluding outliers in asymmetry, curating balanced subsets, or weighting draws toward fairer configurations.
-IV Conclusion
-In this work, we introduced an information–cost framework for quantifying decision complexity in Chess960 starting positions.
-Our analysis shows that the 960 configurations differ markedly in both strategic depth and competitive fairness, challenging the common assumption that all starting positions are equivalent.
-Standard chess, despite centuries of cultural evolution, does not occupy an exceptional location in this landscape: it exhibits a typical initial advantage and moderate total complexity, while displaying above-average asymmetry in decision difficulty.
-This may reflect implicit historical selection for positions that balance richness with accessibility rather than maximal complexity or perfect fairness.
-More broadly, our results demonstrate how concepts from information theory and statistical physics can quantify strategic complexity in deterministic decision-making systems.
-The framework presented here is not specific to chess; it applies to any setting in which optimal choices can be evaluated and the difficulty of distinguishing them can be measured.
-These findings also open perspectives for studying the cultural evolution of board games.
-The fact that the classical starting position is not structurally special raises questions about the ‘phylogenetics’ of chess variants: how the traditional configuration emerged from the broader space of possibilities, and what selection pressures—cognitive, aesthetic, pedagogical, or historical—shaped its adoption.
-Our approach provides quantitative tools for comparing historical games (Shatranj, Xiangqi, Shogi) and for reconstructing the underlying fitness landscape that may have influenced their evolution.
-It also allows one to examine whether successful games cluster in particular regions of complexity space, and whether convergent evolution toward similar complexity profiles occurred independently across cultures.
-Finally, the framework can be extended by incorporating higher-order information-theoretic measures—such as mutual information between successive moves—to probe the dynamical structure of openings.
-Applying similar methods to other randomized-start variants or to entirely different strategic games (Go, Shogi, and modern board games) would test the generality of these ideas and enable systematic cross-game comparisons of strategic complexity.
+II.5 Illustration 1: Searching for jointly balanced positions
+The two fairness metrics introduced above—initial evaluation $E$ and decision asymmetry AA—capture distinct aspects of competitive balance: $E$ measures the expected outcome advantage, while $A$ quantifies the imbalance in cognitive burden between players.
+A natural question is whether these two dimensions are correlated.
+Do positions favoring White in evaluation also impose greater decision complexity on Black?
+Figure 5 plots the asymmetry $A=S_{B}-S_{W}$ against the initial Stockfish evaluation $E$ (depth 15) for all 960 starting positions.
+The correlation is weak ($r\approx 0.15$), indicating that evaluation advantage and decision complexity are largely independent characteristics of a position.
+Positions that are evaluation-balanced do not necessarily yield balanced cognitive demands, and vice versa.
+Figure 5: Complexity asymmetry $A=S_{B}-S_{W}$ versus initial Stockfish evaluation $E$ (Stockfish 17.1, depth 15, single thread) for all 960 Chess960 starting positions.
+The weak correlation ($r\approx 0.15$) indicates near-independence between evaluation balance and decision symmetry.
+The gold star marks position #518 (standard chess).
+The green diamond indicates the configuration that minimizes the normalized distance to the origin in the $(E,A)$ plane at this depth.
+This near-independence motivates the search for configurations that are simultaneously close to zero in both metrics. We therefore define a joint balance score as the normalized distance
+$d=\sqrt{\left(\frac{E}{\sigma_{E}}\right)^{2}+\left(\frac{A}{\sigma_{A}}\right)^{2}},$ (5)
+where $\sigma_{E}$ and $\sigma_{A}$ are the ensemble standard deviations of $E$ and $A$, respectively. This provides a scale-invariant measure of proximity to the ideal balanced point $(E,A)=(0,0)$.
+At depth 15, this criterion identifies position #823 (RKBNQRNB) as the closest configuration to the origin, with $E\approx 0$ pawns and $A\approx+0.26$ bits.
+Importantly, however, the identity of the minimizing position is not universal: as discussed above, outliers and fine rankings can depend on engine depth and evaluation settings.
+The purpose of this construction is therefore not to single out a unique “optimal” starting position, but rather to illustrate how joint fairness criteria can be defined and systematically explored.
+Standard chess (position #518) lies at a moderate distance from the origin, and although standard chess is not extreme in evaluation, it exhibits a non-negligible structural decision asymmetry.
+A deeper chess-specific interpretation of jointly balanced configurations requires further analysis.
+Relevant structural factors may include rook placement (e.g., both rooks on the same flank), king position and castling structure, early tactical constraints, bishop geometry and diagonal accessibility, and the availability of immediate pawn breaks.
+Understanding how such geometric features influence both $E$ and $A$ would provide a bridge between statistical characterization and concrete opening theory.
+II.6 Illustration 2: the most complex positions
+With this consolidated protocol (and depth $15$), the most complex Chess960 initial position is position $524$ (RBNQKNBR), with a total complexity $S_{\mathrm{tot}}=17.17$ bits.
+For comparison, the standard chess starting position (position $518$, RNBQKBNR) has $S_{\mathrm{tot}}=11.20$ bits (and initial evaluation $E=39$ cp).
+Interestingly, this most complex position remains very close to the standard chess configuration, differing only by the permutation of a few pieces (B and N on both king and queen sides).
+Its engine evaluation is also relatively small, $E=24$ cp, which is even closer to equality than the standard chess position.
+This observation confirms that high combinatorial complexity is not associated with a large initial imbalance in the engine evaluation.
+The other configurations with the largest complexity have also values far above that of the standard starting position.
+The five most complex positions identified in our analysis are: position $524$ (RBNQKNBR) with $S_{\mathrm{tot}}=17.166$ ($S_{W}=8.886$, $S_{B}=8.280$); position $444$ (RBNNQKBR) with $S_{\mathrm{tot}}=16.685$ ($S_{W}=8.190$, $S_{B}=8.495$); position $932$ (RBBKRNQN) with $S_{\mathrm{tot}}=15.746$ ($S_{W}=7.383$, $S_{B}=8.362$); position $535$ (RNBKQNRB) with $S_{\mathrm{tot}}=15.540$ ($S_{W}=7.756$, $S_{B}=7.784$); position $372$ (NBBRKRNQ) with $S_{\mathrm{tot}}=15.277$ ($S_{W}=7.896$, $S_{B}=7.381$).
+It would be interesting to understand what these positions have in common and what governs their large complexity values.
+Finally, it is important to stress that these values are obtained at search depth $15$.
+While this depth already provides a reasonable proxy for human-level analysis, deeper searches may modify the ranking of the most complex positions, as discussed above.
+The configurations reported here should therefore be interpreted as belonging to a family of extremely high-complexity starting positions rather than as defining an absolutely fixed ordering across all possible search depths.
+III Discussion and Conclusion
+We introduced an information–cost framework to quantify decision complexity across all 960 Chess960 starting positions.
+Two main results emerge.
+First, the Chess960 ensemble forms a highly heterogeneous landscape: total opening complexity and decision asymmetry vary widely, showing that small rearrangements of back-rank pieces can substantially alter both strategic depth and competitive fairness.
+In particular, the asymmetry $A=S_{B}-S_{W}$ spans several bits, with some positions disproportionately burdening White and others Black.
+A small random sample of starting positions does not, in general, guarantee competitive balance.
+Our framework provides a quantitative basis for designing more principled selection protocols in Freestyle tournaments.
+Second, the classical starting position (position #518) occupies a statistically unremarkable location in this landscape.
+It exhibits a typical evaluation advantage for White, moderate total complexity, and a slightly positive asymmetry indicating a greater cognitive burden for Black.
+Despite centuries of play, the traditional arrangement does not maximize complexity nor minimize imbalance.
+This suggests that its historical persistence may reflect aesthetic symmetry, pedagogical accessibility, or cultural contingency rather than structural optimality.
+More broadly, this work demonstrates how tools from information theory and statistical physics can characterize strategic complexity in deterministic decision systems.
+The framework is not specific to chess: it applies whenever optimal actions can be evaluated and the informational effort required to distinguish them can be quantified.
+Extending this approach to other chess variants, historical games, or different strategic systems (Go, Shogi, and modern board games) would allow systematic comparisons of complexity landscapes and provide new insight into the evolution and design of competitive games.
 Acknowledgements.
 Acknowledgments
-I thank the GothamChess and Blitzstream channels for providing valuable food for thought about chess. I also thank Giordano de Marzo for his suggestion about thinking time and complexity.
+I thank Giordano de Marzo for his suggestion about thinking time and complexity, and Richard Botley for interesting discussions.
+I also thank the GothamChess and Blitzstream channels for providing valuable food for thought about chess.
+Appendix A Appendix A: 960 positions
+he number $960$ follows directly from elementary combinatorics.
+Labeling the back-rank squares (from a to h), the bishops can be placed on any pair of light and dark squares, giving $4\times 4=16$ combinations.
+The queen may then occupy any of the remaining six squares.
+This leaves five squares to be filled by the multiset $\{R,K,R,N,N\}$.
+With 5 squares left, there are $\binom{5}{2}=10$ ways to place the two knights; the remaining three squares must then be occupied by the two rooks and the king, with the king between the rooks, which yields a unique arrangement for these three pieces.
+Altogether, this gives $16\times 6\times 10=960$ possible starting positions.
+This deterministic construction also provides a canonical indexing scheme from 0 to $959$, known as the Scharnagl enumeration algorithm Scharnagl (2004); Wikipedia contributors (2025).
+In this scheme, each starting position is uniquely encoded by three integers: (i) the bishop code, selecting one of the $16$ opposite-colored bishop pairs; (ii) the queen position, selecting one of the $6$ remaining squares after the bishops are placed; and (iii) the N5N code, indexing one of the admissible arrangements of the multiset $\{R,K,R,N,N\}$ on the final five squares under the requirement that the king lies between the two rooks.
+The mapping to an index is given by $\mathrm{idn}\;=\;(\text{bishop code})\;+\;16\times(\text{queen position})\;+\;96\times(\text{N5N code})$.
+For the classical starting position RNBQKBNR, this yields $\mathrm{idn}=6\;+\;16\times 2\;+\;96\times 5=518$ which is the standard numbering for classical chess.
+This indexing provides a compact and reproducible method for scanning all $960$ starting configurations and guarantees a one-to-one correspondence between integers in $[0,959]$ and legal Chess960 starting positions.
+Appendix B Appendix B: Branching factor
+To quantify the complexity evolution across different Chess960 starting positions, we analyze two key metrics: the branching factor $b(n)$ and the position space size $N(n)$ (i.e.
+the number of different positions) as functions of ply depth $n$.
+This analysis characterizes the raw combinatorial complexity of the game tree, independent of strategic considerations or optimal play.
+It quantifies the fundamental decision-making burden at each stage of the game: how many distinct game states a player must consider when planning ahead.
+For each position, we compute $N(n)$ defined as the number of unique legal positions reachable at ply depth $n$.
+We then compute the branching factor $b(n)$ defined as the ratio of consecutive position counts
+$b(n)=\frac{N(n+1)}{N(n)}$ (6)
+This ratio quantifies the average number of accessible successor states per position at depth $n$.
+Figure 6 presents a comparative analysis between position 518 (standard chess) and the ensemble average across all sampled positions.
+Panel (a) shows the evolution of the branching factor $b(n)$ versus ply depth.
+The average branching factor across sampled positions (blue curve with error bars representing one standard deviation) exhibits typical values of order $20$ consistent with previous estimates Barthelemy (2023).
+Note that branching factors around $b\sim 35$, often cited in computer chess literature, correspond to middlegame positions where piece development leads to increased move options Barthelemy (2023).
+Position #518 (green curve) closely tracks this average, with deviations remaining within one standard deviation throughout the analyzed range.
+Panel (b) displays the growth of unique positions $N(n)$ on a semi-logarithmic scale.
+The approximately linear behavior confirms exponential growth, $N(n)\sim\exp(\lambda n)$.
+The extremely small dispersion around the average (smaller than the symbol size) indicates that all starting positions generate state spaces of essentially identical size.
+Note that the effective growth rate $\lambda$ is notably smaller than what would be predicted from the branching factor alone ($\lambda<\ln\langle b\rangle$).
+This reduction arises because different move sequences can converge to the same board configuration—a phenomenon known as transposition in chess terminology—causing the position space to grow more slowly than the space of legal move sequences.
+Figure 6: (a) Average branching factor $\langle b\rangle$ versus ply depth $n$ across all 960 starting configurations (blue), compared with standard chess (position #518, green).
+Error bars indicate one standard deviation.
+The dashed line marks the commonly cited middlegame value $b\approx 35$.
+(b) Number of unique positions $N(n)$ versus ply depth on a semi-logarithmic scale, showing exponential growth $N(n)\sim b_{\mathrm{eff}}^{n}$ with effective branching factor $b_{\mathrm{eff}}\approx 14.41$, significantly smaller than $\langle b\rangle$ due to transpositions.
+From the perspective of branching factors and the number of unique positions, our results show no significant dependence on the initial arrangement; in this sense, all 960 starting positions are equivalent.
+Appendix C Appendix C: $\Delta_{0}$ and time to make a move
+C.1 Effect of $\Delta_{0}$
+The threshold parameter $\Delta_{0}$ controls the definition of a “difficult” move and therefore directly affects both the asymmetry and the total complexity.
+In principle, varying $\Delta_{0}$ should modify the absolute values of these quantities: a smaller $\Delta_{0}$ makes more moves count as difficult, whereas a larger $\Delta_{0}$ selects only the most significant evaluation differences.
+However, our results should not depend critically on the precise choice of $\Delta_{0}$.
+In particular, relative comparisons between starting positions are expected to remain stable provided $\Delta_{0}$ lies within a reasonable range.
+To test this robustness, we computed the asymmetry and total complexity for all 960 starting positions as a function of $\Delta_{0}$, and evaluated the percentile rank of a representative position (the standard position #518) as $\Delta_{0}$ varies.
+Figure 7 shows the evolution of this percentile for engine depth $15$, which is the depth used for most computations in the main text to ensure consistency.
+Figure 7: Percentile rank of the total complexity $S_{\mathrm{tot}}$ and the asymmetry AA for position #518 as a function of $\Delta_{0}$, at engine depth 15.
+The percentile varies smoothly with $\Delta_{0}$, indicating robustness with respect to the choice of threshold.
+We observe that the percentile varies smoothly with $\Delta_{0}$, without abrupt transitions.
+For very small $\Delta_{0}$, almost all moves are classified as difficult; as a consequence, most positions become statistically similar and the percentile of position #518 approaches $50\%$, as expected for a nearly homogeneous ensemble.
+In contrast, for $\Delta_{0}\gtrsim 10$ cp, the percentile stabilizes and varies only weakly with further increases of the threshold.
+These results show that while $\Delta_{0}$ influences absolute complexity values, the relative ranking of positions—and therefore our main conclusions—remain robust over a broad and physically meaningful range of thresholds.
+In particular, the choice $\Delta_{0}=10$ cp (used in the main text for expert-level analysis) lies well within this stable regime.
+C.2 Empirical analysis
+To assess whether the evaluation gap $\Delta$ serves as a meaningful proxy for decision-making complexity, we perform an empirical analysis of players’ thinking times using a large, publicly available database of online games played on Lichess, which includes move-level time stamps Lumbra GigaBase (2024).
+Two successive filters are applied to ensure comparability across games.
+First, we retain only moves played by players with a known Elo rating $\geq 2000$, thereby restricting the analysis to expert-level play and reducing fluctuations associated with lower skill levels.
+Second, we use the PGN TimeControl header (e.g.
+5400+30, 300+3) to classify games by time format.
+We focus here on the Blitz category ($60\leq T_{0}<600$ s, corresponding to 111010 minutes), where time pressure is substantial and thinking time is expected to reflect decision difficulty more directly.
+For this time-control category, we compute the binned average thinking time $\langle\tau\rangle$ as a function of the evaluation gap
+$\Delta=E_{1}-E_{2},$
+where $E_{1}$ and $E_{2}$ denote the engine evaluations (in centipawns) of the best and second-best moves at depth $15$.
+The $\Delta$ axis is partitioned into the following bins: 055, 551010, 10102020, 20205050, 5050100100, 100100200200, 200200500500, and 50050010001000 cp.
+Bins containing fewer than $10$ moves are discarded.
+Error bars represent the standard error of the mean.
+Figure 8: Mean thinking time $\langle\tau\rangle$ as a function of the evaluation gap $\Delta$ for Blitz games with Elo $\geq 2000$.
+Data are computed from $1000$ games (total of $\sim 3{,}300$ moves) extracted from Lumbra GigaBase (2024).
+As shown in Fig.
+8, the average thinking time $\langle\tau\rangle$ decreases clearly with increasing $\Delta$.
+Large values of $\Delta$ correspond to positions in which the best move is clearly superior to the alternatives, resulting in faster decisions.
+Conversely, small $\Delta$ indicates near-indifference between top moves and therefore greater ambiguity, which typically leads to longer reflection times.
+Deviations from strict monotonicity are visible for very small $\Delta$.
+These fluctuations can plausibly be attributed to additional factors, most notably opening preparation: opening positions often exhibit small $\Delta$ (many roughly equivalent moves from the engine’s perspective), yet are played rapidly due to memorized theory.
+Despite these effects, the overall decreasing trend provides strong empirical support for $\Delta$ as a relevant proxy for decision-making complexity.
+Importantly, we observe a broad maximum of the thinking time around $\Delta\approx 10$ cp, suggesting that this scale marks a crossover between positions perceived as genuinely ambiguous and those where the best move becomes sufficiently distinct.
+This empirical observation supports the choice $\Delta_{0}\approx 10$ cp as a natural threshold at expert level.
+C.3 $S(10)$ as a decision-making complexity proxy
+If the entropy $S(10)$ is a meaningful indicator of the intrinsic complexity of a position, one naturally expects it to correlate with the time players spend deciding on a move.
+Intuitively, positions with a large entropy $S(10)$ (i.e., a broad distribution of plausible continuations) should require more cognitive effort, and therefore more time, to resolve.
+To test this hypothesis, we analyze the same freely available database of online games as above (Lichess Lumbra GigaBase (2024)), with move-level time stamps.
+The dataset consists of 1 0001\,000 games, corresponding to a total of 70 15070\,150 individual moves.
+For each position, we compute $S(10)$ using Stockfish at depth 15, and record the time $\tau$ spent by the player on that move.
+Since players of different strengths perceive and handle complexity differently, we account for Elo-dependent effects by adapting the reference threshold $\Delta_{0}$ as a function of the player’s Elo rating.
+Specifically, we use
+$\Delta_{0}=\begin{cases}100,&\text{Elo}<1400,\\[4.0pt] 100-90\,\dfrac{\text{Elo}-1400}{1000},&1400\leq\text{Elo}\leq 2400,\\[4.0pt] 10,&\text{Elo}>2400.\end{cases}$ (7)
+This choice reflects the increasing ability of stronger players to discriminate between close alternatives and to process complex positions more efficiently.
+We then study the dependence of the mean thinking time $\tau$ on $S(10)$, as shown in Fig.
+9.
+As expected, we observe an increase of $\tau$ with $S(10)$.
+Deviations from these trends are visible for very small $\Delta$ or very large $S(10)$.
+These fluctuations can be attributed to additional factors not captured by our complexity measures, such as opening preparation: opening positions often exhibit high complexity and small $\Delta$, yet moves are played rapidly due to memorization.
+Despite these effects, the overall trend provides empirical support for $S(10)$ as a relevant proxy for decision-making complexity in chess.
+It captures, in a quantitative manner, the intuitive notion that positions with many competing continuations or poorly differentiated evaluations are cognitively more demanding and require longer deliberation times.
+Figure 9: Mean thinking time $\tau$ as a function of the entropy $S(10)$, computed with the Elo-dependent threshold $\Delta_{0}$ given by Eq.
+(7).
+Data are obtained from 1 0001\,000 Lichess games (70 15070\,150 moves) Lumbra GigaBase (2024), using Stockfish at depth 15.
+Error bars indicate the standard error of the mean.